### v2

# Context-free graphs and their transition groups

Abstract.
Starting from context-free inverse graphs, we define a new class of groups. We provide closure properties, prove that their co-word problems are context-free, study the torsion elements, and realize them as subgroups of the asynchronous rational group. Furthermore, we use a generalized version of the free product of graphs (see, e.g., [10]) and prove that such a product is context-free inverse closed. We also exhibit an example of a group in our class that is not residually finite and one that is not poly-context-free. These properties make them interesting candidates to disprove both the Lehnert conjecture (which characterizes co-context-free groups as all subgroups of Thompson’s group $V$) and the Brough conjecture (which characterizes finitely generated poly-context-free groups as virtual finitely generated subgroups of direct products of free groups).
1. Introduction
In combinatorial group theory the relationship among formal languages, groups and graphs has been significantly advanced through the investigation of the word problem. Given a group $G$ generated by a finite symmetric set $A$, any element can be represented by a usually non-unique word over $A$. The word problem of $G$ with respect to $A$ is the collection of words representing the identity. Throughout the vast literature, one of the main projects is to fully adapt the famous Chomsky hierarchy to finitely generated groups. The first efforts towards this goal has been made by Anisimov [2], who classified the class of all regular groups as that of all finite groups, and later by Muller and Schupp [22, 23], who have shown that the class of all context-free groups is that of all virtually free groups. Further developments of their renowned theorem were pursued by Ceccherini-Silberstein and Woess [11], and later by Rodaro [24] to generalize it in a more combinatorial setting to certain class of digraphs. We also recall that Herbst has proved that a group is one-counter if and only if it is virtually cyclic (again we have a generalization for this due to Holt, Owens and Thomas ([16]) showing that finitely generated groups having an intersection of finitely many one-counter languages as word problem are virtually abelian) as a intermediate step in the characterization.
The next step of the hierarchy is that of context-sensitive groups. Currently, many intermediate classes have been provided with examples of groups belonging to them: poly-context-free (i.e., its word problem is a finite intersection of context-free languages), co-context-free, co-ET0L, and co-indexed, among others. Several famous examples, such as bounded automata groups and $\mathbb{Z}*\mathbb{Z}^{2}$, were recently exhibited ([6], [1]), in these particular cases they are co-ET0L. On the other hand, the classes of poly-context-free and co-context-free groups come with their universal container conjectures.
Lehnert Conjecture.
A finitely generated group is co-context-free group if and only if it embeds into Thompson $V$.
Here restated in the formulation given by [5].
Brough Conjecture 1.
A finitely generated group is poly-context-free if and only if it is a finitely generated subgroup of a virtual direct products of free groups.
Actually, this second conjecture comes with the following more specific
Brough Conjecture 2.
A finitely generated solvable group is poly-context-free if and only if it is virtually abelian.
Brough Conjecture 3.
If a group $G$ is poly-context-free, then $G$ does not have arbitrarily large finite subgroups.
Our construction heavily relies on the combinatorics of graphs. More precisely, starting from context-free inverse graphs, which are, roughly speaking, deterministic and co-deterministic labeled digraphs with finitely many end-cones, we define a new class of groups called context-free transition groups, which are essentially generated by the permutations of the vertices specified by the labeled directed edges. Despite the nice description of such groups, their possible algebraic structures are yet to be understood, making them interesting on their own. Furthermore, we show that it is, in fact, a subclass of co-context-free groups (Theorem 4.6), and any known poly-context-free group is a context-free transition group (Remark 4.19)). Hence, they fit into the framework of the previous conjectures. In particular, since we provide a non-residually finite transition group (see Section 10), it is worth asking whether the class of context-free transition groups fully contains poly-context-free groups or serves as a tool to construct non-residually finite poly-context-free groups, which would disprove Brough’s Conjecture 1. Note that we also exhibit an example of a transition group that is not poly-context-free (see again Section 10). Moreover, it is well-known that Grigorchuk’s group and $\mathbb{Z}*\mathbb{Z}^{2}$ (see [7]) are two possible counterexamples to Lehnert Conjecture. Even though the first is not a transition group (see Remark 5.7), it would be enlightening to find out whether the second can be realized in this way. On the other hand, in Section 8, we prove that any transition group can be seen as a group of asynchronous bi-lipschitz transducers, though it is not clear whether the construction can be improved to prefix exchange maps.
Sections 2 and 3 are devoted to give some background on inverse graphs and to recall the characterization of context-free inverse graphs. In Section 4 we introduce our class of groups, we show that they are co-context-free and prove some closure properties which allow us to describe any finitely generated subgroup of a virtual direct product of free groups as a transition group. In Section 5 we show some upper bounds of the order of an element with respect to its length, as a consequence we reprove the solvability of the torsion problem, which is already known for the broader class of co-context-free groups. In Section 6 we introduce the subclass of poly-inverse-context-free groups and show some structural properties. In Section 7, we provide structural results on how the transition group is affected when the context-free graph is modified in a small region. In Section 8, we show that such groups can be realized as groups associated with asynchronous transducers, showing that these groups are rational. The main result of Section 9 is that context-free inverse graphs are closed under taking free products of digraphs (Definition 9.1). By exploiting this construction, Section 10 exhibits a non residually finite example of transition group and a transition group containing $\mathbb{Z}^{\infty}$ which, in particular, it means it is not poly-context-free.
2. labeled digraphs, involutive and inverse graphs
Throughout the paper, we assume that $A=\{a_{1}\ldots a_{n}\}$ is a finite set of letters that we call alphabet. The free semigroup on $A$ is denoted by $A^{+}$, and the free monoid on $A$ with empty word $1$ is denoted by $A^{*}$. Any subset of $A^{*}$ is called a language (over $A$). We denote by $F_{A}$ the free group with basis $A$, that is, the group with presentation $\langle A\mid\emptyset\rangle$. If $A^{-1}=\{a^{-1}:a\in A\}$ is the set of formal inverses of the alphabet $A$, we have the involutive closure $A\cup A^{-1}$ with respect to $A$. In this way, from now on we consider our alphabet $A$ to be involutive closed, that is if $a\in A$ also $a^{-1}\in A$, by simply replacing it with its closure. We equip $A^{*}$ with the usual involution $\circ^{-1}:A^{+}\to A^{+}$ defined by $(u^{-1})^{-1}=u$, $(uv)^{-1}=v^{-1}u^{-1}$ for all $u,v\in A^{+}$. A word $w\in A^{*}$ is called (freely) reduced if it does not contain any factor of the form $uu^{-1}$ for some $u\in A^{*}$. It is a well-known fact that given a word $w\in A^{*}$ by successively removing pairs of consecutive inverse letters after a finite number of steps we end up with a unique reduced word $\overline{w}$, called the free reduction of $w$. It is also a well-known fact that the set $R(A)$ of reduced words of $A^{*}$ is a regular language, namely $R=A^{*}\setminus\left(\bigcup_{a\in A}A^{*}aa^{-1}A^{*}\right)$.
Each element of $F_{A}$ may be identified with the equivalence class $[u]=\{w\in A^{*}:\overline{w}=u\}$ for some unique reduced word $u$ which uniquely identifies such class. Thus, the free group $F_{A}$ (with basis $A$) can then be thought of as the set of reduced words in $A^{*}$ with the operation defined by $\overline{u}\cdot\overline{v}=\overline{uv}$. With a slight abuse of notation we will denote the identity of $F_{A}$ by $1$, while for a generic group $G$, we will use the notation $\mathds{1}_{G}$. For a generic subset $L\subseteq A^{*}$, we denote by $\overline{L}=\{\overline{u}:u\in L\}$ the set of all the reduced words that may be obtained from $L$.
An $A$-digraph (or $A$-graph, or just graph) is a tuple $\Gamma=(V,E,\iota,\tau,\lambda)$ where $V=V(\Gamma)$ is the set of vertices, $E=E(\Gamma)$ the set of oriented edges, $A$ the set of labels, $\lambda:E\to A$ is the labeling function, and $\iota:E\to V$, $\tau:E\to V$ are the initial and final (incidence) functions, where for all $e\in E$ we interpret $\iota(e)$, $\tau(e)$ as the initial and terminal vertices of the edge $e$, respectively. Throughout the paper, we mostly adopt a visual approach to such graphs, and we will graphically represent a generic edge $e\in E$ with $\iota(e)=v_{1}$, $\tau(e)=v_{2}$ and $\lambda(e)=a$ as $v_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v_{2}$. Although there are no edges labeled by the empty word, in some arguments sometimes will be useful to consider empty edges of the form $v_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$1$}}}}{{\relbar\joinrel\longrightarrow}}}v_{1}$. We say that $\Gamma$ is deterministic (respectively, complete) if for any vertex $v$ and $a\in A$ there is at most (respectively, at least) one edge $v\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ for some $v^{\prime}\in V(\Gamma)$; and it is co-deterministic if for any vertex $v^{\prime}$ and $a\in A$ there is at most one edge $v\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ incident to $v^{\prime}$. Note that the notion of completeness here comes from automata theory, rather than graph theory, as we are going to view these graphs as language acceptors.
An $A$-digraph $\Gamma^{\prime}=(V^{\prime},E^{\prime},\omega,\tau,\lambda)$ with $V^{\prime}\subseteq V$, $E^{\prime}\subseteq E$ is called a subgraph of $\Gamma$, and we write $\Gamma^{\prime}\subseteq\Gamma$. A walk of length $k$ is a subgraph consisting of a collection of successive edges $p=e_{1}\ldots e_{k}$ with the property $\iota(e_{i+1})=\tau(e_{i})$ for $i=1,\ldots,k-1$. We also sometimes consider infinite one-sided walks $e_{1}e_{2}e_{3}\ldots$, but we remark that infinite one-sided walks may or may not pass through infinitely many vertices. A walk $p$ not containing any subwalk of the form $ee^{-1}$ or any $1$-edge is called reduced. The label of $p$ is the word $\lambda(p)=\lambda(e_{1})\ldots\lambda(e_{k})\in A^{+}$ and if we want to emphasize the label we will sometimes depict such walk as
$p=x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots x_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{k}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{k+1}$
with $e_{i}=x_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a_{i}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i+1}$. The vertices $x_{1},x_{k+1}$ are the initial and terminal vertices of $p$, respectively, and if we do not want to specify them we denote them as $\iota(p),\tau(p)$. For any $x_{i},x_{j}$ with $i<j$, we use $p(x_{i},x_{j})$ to denote the subwalk of $p$ with endpoints $x_{i},x_{j}$. If $\Gamma$ is deterministic given the initial vertex $x_{1}$ and a word $w\in A^{*}$ there exists at most one walk $p$ labeled by $w$ and initial vertex $x_{1}$, which we sometimes denote as $x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y$. In this way, we can also consider the infinite walk $x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, meaning
$x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots y_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y_{k+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots.$
An $A$-digraph is said to be connected if for any two vertices there exists a walk from one to another. In case the walk $p$ has the property $\iota(p)=\tau(p)$, $p$ is called a circuit. When we fix a vertex $x_{0}$ in $\Gamma$, the pair $(\Gamma,x_{0})$ is called a rooted digraph. In this case, we see $(\Gamma,x_{0})$ as a language acceptor by using the distinguished vertex $x_{0}$ as an initial and final state. In this case, the language accepted by $(\Gamma,x_{0})$ is given by the set $L(\Gamma,x_{0})$ formed by the labels of all circuits with initial vertex $x_{0}$. Note that $L(\Gamma,x_{0})$ together with the usual operation of concatenation, is also a submonoid of the free monoid $A^{*}$. More generally, an $A$-automaton has an initial state and a set $F\subseteq V(\Gamma)$ of final states, the language $L(\Gamma,x_{0},F)$ is the set formed by the labels of the walks $p$ with $\iota(p)=x_{0}$, $\tau(p)\in F$. In case $\Gamma$ is finite, $L(\Gamma,x_{0},F)$ belongs to the class of regular languages [18].
The class of $A$-digraphs forms a category where a morphism $\varphi:\Gamma^{(1)}\to\Gamma^{(2)}$ from the digraph $\Gamma^{(1)}=(V_{1},E_{1},\iota_{1},\tau_{1},\lambda_{1})$ to the digraph $\Gamma^{(2)}=(V_{2},E_{2},\iota_{2},\tau_{2},\lambda_{2})$ is a pair of maps $\varphi_{e}:E_{1}\to E_{2}$, $\varphi_{v}:V_{1}\to V_{2}$ that commute with the initial and final maps and with the labeling function:
$\forall g\in E_{1}:\;\;\tau_{2}(\varphi_{e}(g))=\varphi_{v}(\tau_{1}(g)),\,\,\iota_{2}(\varphi_{e}(g))=\varphi_{v}(\iota_{1}(g)),\,\,\lambda_{2}(\varphi_{e}(g))=\lambda_{1}(g)$
Given a subgraph $\Lambda$ of $\Gamma^{(1)}$ we denote by $\psi(\Lambda)$ the subgraph of $\Gamma^{(2)}$ with set of vertices $\varphi_{v}(V(\Lambda))$ and set of edges $\varphi_{e}(E(\Lambda))$. An isomorphism (monomorphism) is a morphism $\varphi$ such that $\varphi_{v}$, $\varphi_{e}$ are both bijective (injective).
Definition 2.1 (involutive and inverse digraphs).
An involutive $A$-digraph is a directed graph $\Gamma$ on the set of labels $A$ with an involution $\circ^{-1}:E\to E$ such that
  • for any $e\in E$, $e^{-1}$ is the opposite edge with $\iota(e^{-1})=\tau(e)$, $\tau(e^{-1})=\iota(e)$;
  • the involution is compatible with the labeling function: $\lambda(e^{-1})=\lambda(e)^{-1}$;
  • no edge is the inverse of itself $e\neq e^{-1}$.
$\Gamma$ is called inverse if in addition to being involutive, it is deterministic and connected, i.e., any two vertices are connected by a walk in $\Gamma$.
In literature, it is common to refer to such a structure as a labeled Serre graph.
Remark 2.2.
In the same way as we observe for alphabet, a digraph can be turn into an involutive one. If $\Gamma$ is a $B$-digraph with $B$ such that $B\cap B^{-1}=\emptyset$, then by adding for each edge $e=x_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}x_{2}$ in $\Gamma$ a new edge $e^{-1}=x_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{1}$ with label on the set of formal inverse $B^{-1}$, we obtain a new involutive digraph on $B\cup B^{-1}$ that denote by $\Gamma^{\pm}$ and called the involutive closure of $\Gamma$.
Note that the involutive closure may be applied to an $A$-digraph with $A$ already involutive closed as usual, in this case, it may happen that we are introducing several “parallel” edges, i.e., edges $e_{1},e_{2}$ so that $e_{1}\neq e_{2}$ with $\iota(e_{1})=\iota(e_{2})$, $\tau(e_{1})=\tau(e_{2})$, $\lambda(e_{1})=\lambda(e_{2})$. Clearly, if $\Gamma$ is an $B$-digraph that is deterministic and co-deterministic, then $\Gamma^{\pm}$ is deterministic. However, if $\Gamma$ is an $A$-digraph that is deterministic and co-deterministic, $\Gamma^{\pm}$ is not in general deterministic because of the parallel edges introduced in the involutive closure.
The notions that we have previously considered for an $A$-digraph, transfer directly to involutive and inverse graphs by adding the compatibility with the involution on the edges. Thus, in the category of involutive (inverse) graphs a morphism $\psi:\Gamma^{(1)}\to\Gamma^{(2)}$ of involutive (inverse) graphs is a morphism of $A$-digraphs with the property that $\psi(e^{-1})=\psi(e)^{-1}$. In particular, considering the inclusion map, the subgraph of an involutive (inverse) graph is also involutive (inverse).
Given a connected involutive graph $\Gamma$, we can equip $\Gamma$ with a structure of a metric space by considering the usual distance given by
$d(u,v)=\min\{n:\mbox{ such that }p=e_{1}\ldots e_{n}\mbox{ is a walk with }\iota(p)=u,\tau(p)=v\}$
A walk $p$ is a geodesic, if $p$ has length exactly $d(\iota(p),\tau(p))$. The subgraph $D_{n}(p)$ consisting of vertices whose length from a fixed vertex $p$ is at most $n$ is called the disk (or ball) of center $p$ and radius $n$; analogously $S_{n}(p)$ denotes the sphere of radius $n$ center at $p$ in the graph. For a connected subgraph $Y\subseteq\Gamma$ the diameter of $Y$ is given by $\delta_{\Gamma}(Y)=\max\{d(y_{1},y_{2}):y_{1},y_{2}\in V(Y)\}$.
The following lemma is a direct consequence of the determinism of inverse graphs.
Lemma 2.3.
Let $\Gamma$ be an inverse $A$-digraph. For any vertex $v\in V(\Gamma)$ and word $u\in A^{*}$ there is at most one walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$ labeled by $u$. Moreover, if there is a walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$, then there is also the (reduced) walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$\overline{u}$}}}}{{\relbar\joinrel\longrightarrow}}}y$.
2.1. Basic facts about morphisms of inverse graphs
The condition of determinism and connectedness of an involutive graph has some consequences when we consider morphisms in the category of inverse graphs. For instance, a morphism $\psi:\Gamma\to\Lambda$ is completely determined by the vertex map $\psi_{v}$. In particular, $\psi$ is a monomorphism if for any $y\in V(\Lambda)$, the preimage $\psi_{v}^{-1}(y)$ contains at most one element and checking if two morphisms are equal is just a matter of verifying if they agree on a single vertex as we will see in the following result.
Lemma 2.4.
Let $\varphi_{1},\varphi_{2}$ be two morphisms from the inverse graph $\Gamma^{(1)}$ to the inverse graph $\Gamma^{(2)}$, then $\varphi_{1}=\varphi_{2}$ if and only if $\varphi_{1}(x)=\varphi_{2}(x)$ for some $x\in V(\Gamma)$.
Proof.
Any walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}y$ is mapped by $\varphi_{1},\varphi_{2}$ into the two walks
$\varphi_{1}(x)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{1}(y)\mbox{ and }\varphi_{2}(x)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi_{2}(y)$
and so by the determinism of $\Gamma^{(2)}$ and condition $\varphi_{1}(x)=\varphi_{2}(x)$ we deduce $\varphi_{1}(y)=\varphi_{2}(y)$. Since $\Gamma^{(1)}$ is connected, we have that $\varphi_{1},\varphi_{2}$ agree on $V(\Gamma^{(1)})$, hence by the determinism of $\Gamma^{(2)}$, they also agree on $E(\Gamma^{(1)})$. ∎
Another interesting feature of inverse graphs, and particularly rooted inverse graphs, is the relationship between accepted languages and morphisms. This connection is contained in the following proposition belonging to folklore, see for instance [24, Proposition 1].
Proposition 2.5.
Let $(\Gamma,x_{0}),(\Lambda,y_{0})$ be two rooted inverse graphs, then $L(\Gamma,x_{0})\subseteq L(\Lambda,y_{0})$ if and only if there is a morphism $\psi:\Gamma\rightarrow\Lambda$ with $\psi(x_{0})=y_{0}$. Moreover, $L(\Gamma,x_{0})=L(\Lambda,y_{0})$ if and only if there is an isomorphism $\psi:\Gamma\rightarrow\Lambda$.
Note that by the previous proposition, we have that, up to isomorphism, there is a unique rooted inverse graph accepting the language $L(\Gamma,x_{0})$. The set of automorphisms of the inverse graph $\Gamma$ will be denoted by $\operatorname{Aut}(\Gamma)$. This is a group acting on $\Gamma$ in a canonical way. In case the quotient graph $\Gamma\setminus\operatorname{Aut}(\Gamma)$ is finite, we say that $\Gamma$ is quasi-transitive. Observe that this last condition is equivalent to the finiteness of $V(\Gamma)\setminus\operatorname{Aut}(\Gamma)$. Any morphism $\varphi:\Gamma\to\Lambda$ between two inverse graphs induces a bijections $\varphi:\operatorname{Star}(\Gamma,v)\to\operatorname{Star}(\Gamma^{\prime},\varphi(v))$ between the star sets in the obvious ways. Thus, morphisms of inverse graphs are immersions of graphs as defined by Stallings [25]. In particular, if the inverse graphs are complete, by Lemma 2.3 every morphism between two complete inverse graphs is a covering, that is, an immersion of graphs $f:\Gamma\to\Lambda$ so that $f$ maps the sets $V(\Gamma)$ and $E(\Gamma)$ respectively onto $V(\Lambda)$ and $E(\Lambda)$. One central feature of coverings is the possibility of lifting walks, this is contained in the following characterization.
Proposition 2.6 (Proposition 2.3 of [15]).
Let $\varphi:\Gamma^{\prime}\to\Gamma$ be a morphism of inverse graphs. Then, for each vertex $v\in V(\Gamma)$ and vertex $v^{\prime}\in\varphi^{-1}(v)$, every walk $p$ in $\Gamma$ with $\iota(p)=v$ lifts to a unique walk $p^{\prime}$ starting at $v^{\prime}$, i.e., $\varphi(p^{\prime})=p$.
3. Context-free grammars, pushdown automata, and configurations graphs
There are several ways to define a context-free language, either via the notion of grammar, or using the notion of pushdown automaton. For the sake of completeness, we briefly recall the main notions that are used in the paper, and we refer the reader to [18] for further details.
Definition 3.1.
A context-free grammar is a tuple $\mathcal{G}=(V,\Sigma,P,S)$ where $V$ is a finite set of non-terminal symbols (or variables), $\Sigma$ is the finite alphabet of terminal symbols that is disjoint from $V$, $S\in V$ is the starting symbol, and $P\subseteq V\times(V\cup\Sigma)^{*}$ is a finite set called productions.
Usually production are denoted by $X\Rightarrow_{\mathcal{G}}\alpha$, $\alpha\in(V\cup\Sigma)^{*},X\in V$. We extend $\Rightarrow_{\mathcal{G}}$ by putting $\alpha X\gamma\Rightarrow_{\mathcal{G}}\alpha\beta\gamma$ whenever $X\Rightarrow_{\mathcal{G}}\beta\in P$, and we denote by $\Rightarrow_{\mathcal{G}}^{*}$ the reflexive and transitive closure of $\Rightarrow_{\mathcal{G}}$. The language generated by the grammas $\mathcal{G}$ is $L(S)=\{u\in\Sigma^{*}:S\Rightarrow^{*}_{\mathcal{G}}u\}$, and in this case we say that it is a context-free language.
[…]
In what follows, we will concentrate our attention on deterministic pushdown automata that are real-time, so without $1$-moves. Pushdown automata without $1$-moves have configurations graphs without edges labeled with the empty word $1$, so it is the natural framework when considering configurations graphs that are inverse $\Sigma$-digraphs. It is clear that to have $\Gamma(M)$ that is a deterministic $\Sigma$-digraph, we need to assume $M$ deterministic. The natural question is when $\Gamma(M)$ is also co-deterministic, since in this case we may transform $\Gamma(M)$ into an inverse graph $\Gamma(M)^{\pm}$ by simply adding for every edge $(q,\xi x)\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}(p,\xi\gamma)$ the reverse one $(p,\xi\gamma)\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}(q,\xi x)$ (see Remark 2.2). In particular, it is of interest to see when $\Gamma(M)^{\pm}$ is the configurations graph of some deterministic PDA. We shall see that this happens if $M$ has a reverse transition function, so $M$ is a reversible pushdown automaton. A reversible pushdown automaton $M=(Q,\Sigma,\Theta,\delta,q_{0},\perp,F)$ is characterized by having a reverse transition function $\delta_{R}:Q\times\Sigma\times\Theta\to Q\times\Theta^{*}$ such that is able to locally reverse the computation, i.e.,
[…]
We remark that the initial assumption that $M$ is real-time is not strict from the language point of view since by [19, Theorem 5] for every reversible pushdown automaton $N$ there is a real-time reversible PDA $M$ such that $L(M)=L(N)$. To deal with inverse graphs, a new class of PDAs that is more suitable to consider languages on involutive closed alphabet $\Sigma\cup\Sigma^{-1}$ in [24] has been introduced.
Definition 3.2 (Inverse PDA and inverse-context-free languages).
A real-time deterministic PDA $M=(Q,\Sigma\cup\Sigma^{-1},\Theta,\delta,q_{0},\perp,F)$ is called inverse if $\delta$ may be partitioned into two functions: $\delta^{+}$ which is the restriction of $\delta$ to $Q\times\Sigma\times\Theta$ and $\delta^{-}$ which is the restriction of $\delta$ to $Q\times\Sigma^{-1}\times\Theta$ such that $\delta^{+}$ is the reverse transition of $\delta^{-}$ and vice versa in the following sense:
[…]
A language $L\subseteq(\Sigma\cup\Sigma^{-1})^{*}$ is called inverse-context-free if $L=L(M)$ for some inverse PDA $M$ on $\Sigma\cup\Sigma^{-1}$.
It is clear from the definition and the fact that $\delta^{+},\delta^{-}$ are functions, that the configurations graph $\Gamma(M)$ is inverse if and only if $M$ is an inverse PDA. For an invertible PDA $M$ let us set $M^{+}=(Q,\Sigma,\Theta,\delta^{+},q_{0},\perp,F)$. It is straightforward to see that a PDA $M$ is inverse if and only if $M^{+}$ is reversible.
3.1. Inverse context-free graphs
The machine point of view that we have seen before can be reformulated into a more graph-theoretic setting via the notion of context-free graph. The starting point is the notion of context-free group given in [22]. In their paper, Muller and Schupp called a finitely generated group $G$ presented by $\langle A\mid\mathcal{R}\rangle$ context-free if the word problem $WP(G;A)$ is a context-free language. With our notation, this language coincides with the set $L(\operatorname{Cay}(G;A),\mathds{1}_{G})$ of words labeling a circuit starting and ending at the group identity.
Definition 3.3.
A rooted inverse graph $(\Gamma,x_{0})$ is called context-free if $L(\Gamma,x_{0})$ is a context-free language.
This definition is equivalent to a more geometric one that involves the notion of end-cones. We briefly recall it. Henceforth, we denote by $d(x,y)$ the usual distance in the inverse graph $\Gamma$, and put $\|v\|_{x_{0}}=d(x_{0},v)$. In [23] Muller and Schupp introduced the central notion of end-cone together with the idea of end-isomorphism of such subgraphs. For a vertex $v$ with $\|v\|_{x_{0}}=n$, the end-cone at $v$ is the (inverse) subgraph $\Gamma(v,x_{0})$ that is the connected component of $\Gamma\setminus D_{n-1}(x_{0})$ containing $v$, and we denote by $\Delta(v,x_{0})=\{y\in\Gamma(v,x_{0}):\|y\|_{x_{0}}=n\}$ the set of frontier vertices (or vertices) of $\Gamma(v,x_{0})$. An end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma(w,x_{0})$ between the end-cones $\Gamma(v,x_{0}),\Gamma(w,x_{0})$ is an isomorphism of inverse graphs with the property of preserving their frontier vertices: $\psi(\Delta(v,x_{0}))=\Delta(w,x_{0})$. We sometimes write $\Gamma(v,x_{0})\simeq_{end}\Gamma(w,x_{0})$ to emphasize the existence of an end-isomorphism between the two end-cones.
Definition 3.4.
The rooted graph $(\Gamma,x_{0})$ is called a context-free graph in the sense of Muller-Schupp if up to end-isomorphism there are finitely many end-cones, i.e., there are inverse graphs $\Gamma_{1},\ldots,\Gamma_{N}$ such that for any end-cone $\Gamma(v,x_{0})$ of $\Gamma$ there is an end-isomorphism $\psi:\Gamma(v,x_{0})\to\Gamma_{j}$ for some $j\in[1,N]$.
In [23] it is shown that the graph $(\Gamma(M),x_{0})$ of the instantaneous descriptions of a pushdown automaton is a context-free graph, and conversely given a context-free graph $(\Gamma,x_{0})$ there is a pushdown automaton $M$ such that $(\Gamma(M),x_{0})$ is isomorphic to $(\Gamma,x_{0})$. The following characterization relates all the notions that we have seen so far.
Proposition 3.5 (Proposition 7 of [24]).
For an inverse graph $\Gamma$ the following conditions are equivalent:
  1. (1)
    for any non-empty finite set $F\subseteq V(\Gamma)$ the language $L(\Gamma,x_{0},F)$ is context-free;
  2. (2)
    the language $L(\Gamma,x_{0})$ is context-free, hence $(\Gamma,x_{0})$ is context-free;
  3. (3)
    the rooted graph $(\Gamma,x_{0})$ is context-free in the sense of Muller-Schupp;
  4. (4)
    there exists an inverse PDA $M$ with initial state $q_{0}$ and initial stack symbol $\perp$ such that the rooted configurations graph $(\Gamma(M),(q_{0},\perp))$ is isomorphic to $(\Gamma,x_{0})$;
  5. (5)
    for any vertex $y_{0}\in V(\Gamma)$, the rooted graph $(\Gamma,y_{0})$ is context-free in the sense of Muller-Schupp.
4. Transition groups of inverse graphs and co-context-free groups
In this section, we introduce a new subclass of co-context-free groups. For an $A$-digraph $\Gamma$, the set of everywhere-circuits is
(1) $\mathcal{L}(\Gamma)=\{u\in A^{*}:\forall p\in V(\Gamma)\,\mbox{ there is a circuit }p\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ in }\Gamma\}=\bigcap_{p\in V(\Gamma)}L(\Gamma,p)$
It is easy to see that if $\Gamma$ is an inverse $A$-digraph that is also complete, then $\mathcal{L}(\Gamma)$ is a normal subgroup of $F_{A}$. In this setting, we give the following definition.
Definition 4.1.
For a complete inverse $A$-digraph $\Gamma$, each letter $a\in A$ defines a permutation on the set of vertices $a:V(\Gamma)\to V(\Gamma)$. The transition group $\mathcal{G}(\Gamma)$ of the inverse graph $\Gamma$ is the group generated by composing all these permutations.
It is easy to see that $\mathcal{G}(\Gamma)$ may be equivalently obtained by considering the quotient of the free group $F_{A}$ generated by $A$ by the normal subgroup $\mathcal{L}(\Gamma)$. There is a natural right action of $\mathcal{G}(\Gamma)$ on the set of vertices $V(\Gamma)$: for any $g\in\mathcal{G}(\Gamma)$ and vertex $v\in V(\Gamma)$, $v\cdot g\in V(\Gamma)$ is the vertex for which there is a walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}(v\cdot g)$ in $\Gamma$ for any $w\in A^{*}$ that represents $g$. We sometimes also put $v\cdot w=v\cdot g$ for any $w\in A^{*}$ representing $g$ in $\mathcal{G}(\Gamma)$.
Example 4.2.
By taking $\Gamma$ either as a finite $n$-cycle or an infinite line and by suitably labeling all edges either by $a$ or $a^{-1}$ from a single-letter alphabet $A=\{a\}$, we can turn $\Gamma$ into a complete inverse graph so that its transition group is readily seen to be either isomorphic to $\mathbb{Z}/n\mathbb{Z}$ or $\mathbb{Z}$, so that all these groups are transition groups.
Remark 4.3.
The transition group is the syntactic monoid of the language $L=L(\Gamma,x_{0})$ (where $x_{0}$ is any chosen vertex of $\Gamma$). Indeed, it is not difficult to see that if one considers the syntactic congruence $\sim_{\Gamma}$ on $A^{*}$ defined by $u\sim_{L}v$ if for all $x,y\in A^{*}$ $xuy\in L$ if and only if $xvy\in L$, then since $\Gamma$ is inverse, we have that $A^{*}/_{\sim_{L}}$ is isomorphic to $\mathcal{G}(\Gamma)$. If we remove the completeness condition, the syntactic monoid is an inverse monoid. For a general introduction on syntactic monoids of a formal language see [13].
Note that the rooted graph $(\Gamma,p)$ may be naturally identified with the Schreier graph of the stabilizer subgroup $H=\mathcal{G}(\Gamma)_{p}=\{g\in\mathcal{G}(\Gamma):g(p)=p\}$. This fact follows from Proposition 2.5 and the equality $L(\Gamma,p)=L(\operatorname{Sch}(H,A),H)$.
We actually want to focus on a slightly more general framework: we consider a finite collection of inverse complete $A$-digraphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ and we see that, as before, we can consider the permutations induced by the letters of $A$. In this way, we speak about the transition group
$\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)}):=F_{A}/_{\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})}$
with $\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})$.
It will be useful to see a transition group over a finite union of graphs as a the subgroup of a direct product of transition groups over a single graph. Namely, we have the following lemma.
Lemma 4.4.
If $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ are inverse $A$-graphs that are complete, then the transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is isomorphic to a subgroup of $\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$.
Proof.
Let $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ and consider the map $\psi:G\to\mathcal{G}(\Gamma^{(1)})\times\ldots\times\mathcal{G}(\Gamma^{(k)})$ sending an element $g\in G$ represented by a word $u\in A^{*}$ to the corresponding element $(\pi_{1}(u),\ldots,\pi_{k}(u))$, where $\pi_{i}:A\to\mathcal{G}(\Gamma^{(i)})$ are the natural projection maps. Now by definition
$\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})=WP(G;A).$
Hence $\psi$ is a well defined homomorphism since an element $g=\mathds{1}$ in $G$ is equivalent to $u\in WP(G;A)$, so $u\in\mathcal{L}(\Gamma^{(i)})$ and thus $\pi_{i}(u)=\mathds{1}$ for all $i=1,\ldots,k$. Moreover, $\psi$ is injective since $\pi_{i}(u)=\mathds{1}$ for all $i=1,\ldots,k$ is equivalent to $u\in\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=WP(G;A)$. ∎
Before proving the next lemma we need to recall the following definition. For a language $L\subseteq A^{*}$, the shift of $L$, is the language $S_{A}(L)=\{xy\in A^{*}:yx\in L\}$ formed by all the words that may be obtained by a cyclic shift of some word of $L$. If it is clear from the context, we drop the subscript $A$ and we simply refer to $S(L)$. It is a well-known fact (see for instance Ex. 6.4 of [18]) that for a context-free language $L$, $S(L)$ is also context-free.
Lemma 4.5.
Let $\mathcal{G}(\Gamma)$ be transition group of a complete context-free graph $\Gamma$ and let $g$ be a non trivial element of $\mathcal{G}(\Gamma)$. There exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$ and $v\cdot g\neq v$.
Proof.
Let $q\in V(\Gamma)$ be the vertex from which there is a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}p$ with $p\neq q$ that is not a circuit. Consider the infinite one-sided walk $\pi=q\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, which exists since $\Gamma$ is complete. In $\pi$ choose a vertex $v$ belonging to the frontier vertices $\Delta(v,x_{0})$ such that $\pi\subseteq\Gamma(v,x_{0})$ (it exists since every vertex belongs to some set of frontier vertices and $\pi\subseteq\Gamma=\Gamma(x_{0},x_{0})$). Equivalently, $v$ is chosen among the vertices $\pi$ whose distance from $x_{0}$ is minimum. Using the vertex $v$ we may find a cyclic-shift $yx$ of $xy=g$ such that $\pi$ factors as follows:
$\pi=q\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}p\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{2}\smash{\stackrel{{\scriptstyle\text{\tiny{$(yx)$}}}}{{\relbar\joinrel\longrightarrow}}}z_{3}\ldots z_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$(yx)^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$
where $v=z_{k}$ for some $k\geq 1$. We claim that there is a walk $v=z_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$yx$}}}}{{\relbar\joinrel\longrightarrow}}}z_{k+1}$ in $\Gamma(v,x_{0})$ that is not a circuit. Indeed, if $z_{k}=z_{k+1}$, then by the determinism of $\Gamma$ we would have $q=p$, a contradiction. ∎
For the next theorem we need to know that the union of finitely many context-free languages is context-free, and for a deterministic context-free language the complement is also deterministic context-free, see [18].
Theorem 4.6.
Let $\Gamma$ be a context-free inverse $A$-digraph that is also complete, then $\mathcal{L}(\Gamma)\in{\bf CoCF}$.
Proof.
Let $\{\Gamma_{1},\ldots,\Gamma_{N}\}$ be the end-cones types of $\Gamma$ and let $\Delta_{i}$ be the corresponding frontier vertices. For each $q\in\Delta_{i}$ we denote by
$N(q)=\{u\in A^{*}:q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\mbox{ is a walk in }\Gamma_{i}\mbox{ with }p\neq q\}$
the set of words that do not label a circuit at $q$ in $\Gamma_{i}$. We claim:
$\mathcal{L}(\Gamma)^{c}=\bigcup_{i=1}^{N}\left(\bigcup_{q\in\Delta_{i}}S\left(N(q)\right)\right)$
Let us prove that the right hand side is contained in the left hand side. Note that every walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Gamma_{i}$ that does not label a circuit at $q$, has the same property in $\Gamma$ in any subgraph that is end-isomorphic to $\Gamma_{i}$, i.e., $u\in\mathcal{L}(\Gamma)^{c}$. Moreover, take the word $yx$ such that $xy=u\in N(q)$, and factor a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ with $p\neq q$ in $\Gamma$ as $q\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}p$. If $yx$ labeled a circuit at $v$, then by the determinism of $\Gamma$ we would have $p=q$, a contradiction. Hence, $yx\in S(N(q))$ does not label a circuit at $v$ in $\Gamma$, i.e., $yx\in\mathcal{L}(\Gamma)^{c}$.
Let us now prove the other inclusion. So let $g\in\mathcal{L}(\Gamma)^{c}$, by Lemma 4.5 there exists an end-cone $\Gamma(v,x_{0})$ s.t. the infinite walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ with $\mathring{g}\in S(\{g\})$ and $v\cdot g\neq v$. Hence, $g\in S(N(w))$ where $w$ is the frontier vertex in the end-cone type $\Gamma_{j}$ that is end-isomorphic to the end-cone $\Gamma(v,x_{0})$, and this concludes the proof of the other inclusion. To show that $\bigcup_{i=1}^{N}\left(\bigcup_{q\in\Delta_{i}}S\left(N(q)\right)\right)$ is context-free, it is enough to show that each $N(q)$ is context-free, since context-free languages are closed under the cyclic-shift operator, and a finite union of context-free languages is context-free. Now, for any $q\in\Delta_{i}$, $N(q)=L(\Gamma_{i},q)^{c}$ and since $\Gamma_{i}$ is, in particular, a context-free graph, we have that $L(\Gamma_{i},q)$ is a deterministic context-free language since by Proposition 3.5, $\Gamma_{i}$ may be seen as the configurations graph of an inverse pushdown automaton accepting $L(\Gamma_{i},q)$. Now, since deterministic context-free languages are closed under complementation, we conclude that $L(\Gamma_{i},q)^{c}$ is also context-free, and this concludes the proof of the proposition. ∎
Remark 4.7.
In the previous proof we have shown that each language $N(q)$ is a deterministic context-free language. If $S(N(q))$ were deterministic for all $q\in\Delta_{i}$, then, by taking the complement, we would deduce that $\mathcal{L}(\Gamma)$ is an intersection of finitely many context-free languages, i.e., $\mathcal{L}(\Gamma)$ is poly-context-free. However, it seems that the cyclic closure of a context-free language is generally far from being deterministic, making the class of transition groups of context-free inverse graphs quite different from that of poly-context-free groups. In fact, we are going to find an example of a transition group which is not poly-context-free (see Definition 10.4).
As an immediate consequence of the previous proposition, we obtain the following theorem.
Theorem 4.8.
Let $\Gamma$ be a context-free inverse $A$-digraph that is also complete. Then, the transition group $\mathcal{G}(\Gamma)$ is CoCF.
In the theorem above, the connectedness of $\Gamma$ is not strictly necessary and since CoCF languages are closed under finite intersections we have
Theorem 4.9.
If $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ are context-free inverse $A$-graphs that are complete, then, the transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is CoCF.
Definition 4.10 (CF-TR groups).
In view of Theorem 4.9, we may define the subclass of CoCF groups formed by all the transition groups of a finite disjoint collection of context-free inverse $A$-digraphs, denoted by CF-TR.
Question 4.11.
There are several natural questions for such a class of groups.
  1. (1)
    What is the relationship with the class of poly-context-free groups?
  2. (2)
    Is the class CF-TR strictly included into CoCF?
There is a natural characterization of such groups in terms of certain coverings. We say that a covering $\psi:\Lambda\to\Gamma$ is transition-preserving if for any for any walk $q^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ with $q^{\prime}\neq p^{\prime}$ in $\Lambda$ there is a walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p$ in $\Lambda$ such that the corresponding walk $\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(p)$ in $\Gamma$ is not a circuit. We have the following fact.
Proposition 4.12.
If $\psi:\Lambda\to\Gamma$ is a transition-preserving covering between two complete $A$-digraphs, then $\mathcal{G}(\Gamma)\simeq\mathcal{G}(\Lambda)$. In particular, a group $G$ is CF-TR if and only if there is a transition-preserving covering of the Cayley graph $\operatorname{Cay}(G;A)$ onto a context-free $A$-digraph.
Proof.
Consider the map $\phi:\mathcal{G}(\Lambda)\to\mathcal{G}(\Gamma)$ defined as follows: for any $g\in\mathcal{G}(\Lambda)$ and $u\in A^{*}$ which represents $g$, we define $\phi(g)=h$ where $h\in\mathcal{G}(\Gamma)$ is the element defined by $u$. It is a well-defined map since if $v$ represents $g$ too, then $uv^{-1}\in\mathcal{L}(\Lambda)$ labels a circuit everywhere in $\Lambda$, and so also in $\Gamma$ since $\psi$ is a covering. Hence, $v$ represents $h$ as well. It is routine to check that it is a morphism of groups which is also surjective. Let us prove that $\phi$ is injective. Let $u\in\mathcal{L}(\Gamma)$, by the path lifting property (Proposition 2.6), any circuit $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}v$ in $\Gamma$ lifts to a walk $v^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ with $\psi(v^{\prime})=\psi(v^{\prime\prime})=v$. Now, if $u\notin\mathcal{L}(\Lambda)$, by the transition-preserving property we would have that $u\notin\mathcal{L}(\Gamma)$, a contradiction.
Let us prove the characterization. It is straightforward to check that the transition group of the Cayley graph $\operatorname{Cay}(G;A)$ is isomorphic to $G$. Thus, if there is a covering $\psi:\operatorname{Cay}(G;A)\to\Gamma$ onto a context-free graph $\Gamma$, then by the first part of this proposition $\mathcal{G}(\Gamma)\simeq G$. Conversely, if $G=\mathcal{G}(\Gamma)$ and we fix a root $x_{0}$ in $\Gamma$, then the morphism $\psi:\operatorname{Cay}(G;A)\to\Gamma$ defined on the vertices by $\psi(v)=y$ whenever $1\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}v$ is a walk in $\operatorname{Cay}(G;A)$ and $y=x_{0}\cdot g$ and extended to a morphism of inverse $A$-digraphs, is a covering that is also transition-preserving. ∎
This class is closed under taking finitely generated subgroups, direct products and finite extensions. In order to prove that, we need some preliminary lemma.
We fix $H$ a subgroup of $G$ that is finitely generated by $B\subseteq A^{*}$ and we set $n:=\max_{B}\ell(b)$ with $\ell(b)$ the length in terms of $A$.
Lemma 4.13.
Let $g\in H$ be a non-trivial element. Then, there exist an end-cone $\Gamma(v,x_{0})$ and a vertex $\widetilde{v}\in\Gamma(v,x_{0})$ with $d(v,\widetilde{v})\leq n$ such that the infinite walk $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$ for some $\mathring{g}\in S_{B}(\{g\})$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$.
Proof.
By Lemma 4.5 there exists an end-cone $\Gamma(v,x_{0})$ such that the infinite walk $\pi=x\smash{\stackrel{{\scriptstyle\text{\tiny{$\widehat{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(v,x_{0})$. Now, by definition of $n$, up to removing a walk of at most $n$ edges to $\pi$, we obtain a walk $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ with $d(v,\widetilde{v})\leq n$ and $\mathring{g}\in S_{B}(\{g\})$. ∎
Proposition 4.14.
The class of CF-TR is closed by taking finitely generated subgroups.
Proof.
Suppose first that we are dealing with one complete context-free graph, so that $H$ is a finitely generated subgroup of $\mathcal{G}(\Gamma)$, where $\Gamma$ is an inverse graph that is also context-free. Let $B\subseteq A^{*}$ be the involutive closed generating set and let $\Delta_{1},\ldots,\Delta_{N}$ be the frontier points of the end-cone types $\{\Gamma_{1},\ldots,\Gamma_{N}\}$. We set $\Delta_{i}^{n}:=\{v\in\Gamma_{i}\mid d(x,v)\leq n\mbox{ for some }x\in\Delta_{i}\}$. With a slight abuse of notation we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We do the same for the set $\Delta_{i}^{n}$. By definition of transition group, we need to show that $H$ is the transition group of a finite union of complete context-free $B$-digraphs $\Theta_{B}$. We claim
$\Theta_{B}=\bigcup_{i=1}^{N}\bigcup_{v\in\Delta_{i}^{n}}(\Gamma^{B},v)$
where $(\Gamma^{B},v)$ is given by the vertices $y$ of $\Gamma$ such that there exists $w\in B^{*}$ and $v\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y$. The edges are $y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}y_{2}$ whenever $w\in B$ and there exists a walk labeled by $w$ from $y_{1}$ to $y_{2}$ in $\Gamma$. It is not difficult to see that since $B$ is involutive closed and $\Gamma$ is inverse, then also $(\Gamma^{B},v)$ is inverse. Moreover, it is complete and context-free.
Let us show that the complement of the word problem of $H$ and $\mathcal{G}(\Theta_{B})$ coincide, as this will of course imply that their word problems coincide. Let $g\in G(\Theta_{B})$ be a non trivial element, then there exists $\widetilde{v}\in(\Gamma^{B},v)$ for some $v\in\Delta_{i}^{n}$ such that $\widetilde{v}\cdot g\neq\widetilde{v}$. It follows that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime}$ with $\widetilde{v}\neq v^{\prime}$ is also a walk in $\Gamma$ and so $g\neq\mathds{1}_{H}$ in $H$.
Conversely, suppose that $\mathds{1}_{H}\neq g\in H$, by Lemma 4.13 there exists an end-cone $\Gamma(x,x_{0})$ and a vertex $v\in\Gamma(x,x_{0})$ with $d(x,v)\leq n$ such that $v\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(x,x_{0})$ and $v\cdot\mathring{g}\neq v$ for some $\mathring{g}\in S_{B}(\{g\})$. Now $\Gamma(x,x_{0})$ is end-isomorphic to some $\Gamma_{i}$, hence we can find $\widetilde{v}\in\Delta_{i}^{n}$ such that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma_{i}$ and $\widetilde{v}\cdot\mathring{g}\neq\widetilde{v}$. A fortiori, $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq(\Gamma,\widetilde{v})$, whence $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subset(\Gamma^{B},\widetilde{v})$. Now in $(\Gamma^{B},\widetilde{v})$, we can read $g$ on a walk $v^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$s$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ such that $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$s$}}}}{{\relbar\joinrel\longrightarrow}}}v^{\prime\prime}$ is a subwalk of $\widetilde{v}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{g}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ and where $v^{\prime}\neq v^{\prime\prime}$, otherwise determinism would imply $\widetilde{v}\cdot\mathring{g}=\widetilde{v}$. In particular, $g$ is non trivial in $G(\Theta_{B})$.
The previous argument may easily be adapted when considering a finite family of disjoint complete context-free inverse graphs. ∎
Note that the previous proposition gives a constructive way to compute context-free inverse graphs defining a finitely generated subgroup of a transition group of a context-free graphs $\Gamma$.
Proposition 4.15.
The class of CF-TR is closed by taking direct products.
Proof.
Let $\mathcal{G}(\Gamma_{i})$ be a transition group of a context-free inverse $A_{i}$-digraph with $i=1,2$. We can assume that $A_{1}\cap A_{2}$ is empty, otherwise we change accordingly the labels in $A_{2}$. Now we consider $\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2}$ where $\widehat{\Gamma}_{i}$ is a $A_{1}\cup A_{2}$-digraph in which for every vertex there are $|A_{j}|$ loops each labeled by an element in $A_{j}$ with $j\neq i$. It is straightforward that $\mathcal{G}(\widehat{\Gamma}_{1}\sqcup\widehat{\Gamma}_{2})$ is the desired direct product. ∎
The technique to prove the last closure property is inspired by [17] and, in particular, it is based on the inflated graph construction introduced in [24] combined to the strategy in the proof of Proposition 4.14.
Definition 4.16.
Let $H$ be a finitely generated group which is a finite index subgroup of $G$ and let $T$ be a right transversal for $G$ with respect to $H$ containing the identity. For each $y\in A\cup T-\{\mathds{1}_{G}\}$ and $t\in T$, we fix a word $h_{t,y}\in A^{*}$ such that $ty=_{G}h_{t,y}\widetilde{t}$ for some $\widetilde{t}\in T$. We define the Schreier automaton $\mathcal{A}$ over $A\cup T$ to be the finite state automaton with set of states $T$ , initial state $\mathds{1}_{G}\in T$ and $t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid h_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}$ when $ty=_{G}h_{t,y}\widetilde{t}$.
Starting from a Schreier automaton, we can build one that is inverse, i.e.
$t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid h_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{t}\ \Longleftrightarrow\widetilde{t}\ \smash{\stackrel{{\scriptstyle\text{\tiny{$y^{-1}\mid h^{-1}_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}t,$
without changing the groups involved $H$ and $G$. The way we do so is by enlarging the alphabet for the labels from $A\cup T$ to $A\cup\{T,T^{-1}\}$ (remember $A$ is already inverse by assumption) and by adding the inverse transitions $\widetilde{t}\ \smash{\stackrel{{\scriptstyle\text{\tiny{$y^{-1}\mid h^{-1}_{t,y}$}}}}{{\relbar\joinrel\longrightarrow}}}t$. The fact that we are still dealing with $H$ and $G$ follows from the equation $\widetilde{t}y^{-1}=_{G}h^{-1}_{t,y}t$, which is true if and only if $ty=_{G}h_{t,y}\widetilde{t}$. We call it the inverse Schreier automaton and with a slight abuse of notation we denote it again by $\mathcal{A}$.
Definition 4.17.
Let $\Gamma$ be a complete inverse $A$-digraph and let $H$ be the associated transition group. Consider $G$ to be a group with $H$ as a finite index subgroup and $\mathcal{A}$ the corresponding Schreier automaton over $A\cup T$. The inflated graph $\mathcal{A}\ltimes(\Gamma,x)$ with $x\in V(\Gamma)$ is the rooted $A\cup T$-digraph defined as follows: the set of vertices is $T\times V(\Gamma)$, a generic edge is $(t,p)\smash{\stackrel{{\scriptstyle\text{\tiny{$y$}}}}{{\relbar\joinrel\longrightarrow}}}(t^{\prime},p^{\prime})$ with $t\smash{\stackrel{{\scriptstyle\text{\tiny{$y\mid u$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\in\mathcal{A}$ and $p\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}$ a walk in $\Gamma$ and the root is $(\mathds{1}_{G},x)$.
It is straightforward that $\mathcal{A}\ltimes\Gamma$ is complete and if $\Gamma$ is context-free, so is $\mathcal{A}\ltimes\Gamma$ (see the proof of [24, Lemma 11]).
Proposition 4.18.
The class of CF-TR is closed by taking finite index overgroups.
Proof.
Let $G$ be a finitely generated group and $H=\mathcal{G}(\Gamma)$ a finite index subgroup of $G$. Let $T$ be a transversal of $G$ with respect to $H$ containing the identity, so that $H=\langle A\rangle$ and $G=\langle A\cup T\rangle$, and let $\mathcal{A}$ be the corresponding Schreier automaton.
Let $\{\Delta_{1},\ldots,\Delta_{N}\}$ be the frontier points, and arguing as we did in Proposition 4.14 with a slight abuse of notation, we identify the frontier vertices of $\Delta_{i}$ with the frontier vertices of an end cone $\Gamma(v,x_{0})$ of $\Gamma$ that is end-isomorphic to $\Gamma_{i}$. We claim that $G$ is the transition group of
$\Theta_{\mathcal{A}}:=\bigcup_{i=1}^{N}\bigcup_{v\in\Delta_{i}}(\mathcal{A}\ltimes(\Gamma,v),(v,\mathds{1}_{G})).$
All we need to prove is that $g\neq 1$ in $G$ if and only if there exists $i\in\{1,\ldots,N\}$, $v\in\Delta_{i}$ and $q\in\mathcal{A}\ltimes(\Gamma,v)$ s.t. $q\cdot g\neq q$. Indeed the latter is exactly the requirement for an element to be non trivial in $\mathcal{G}(\Theta_{\mathcal{A}})$.
First of all, we notice that $g$ is non trivial in $G\setminus T$ if and only if at least one of the following occurs
  1. (1)
    there exists $t\in T$ s.t. $tg=_{G}ht^{\prime}$ for some $t^{\prime}\neq t$ and $h\in\{h_{t,y}\}_{t\in T,y\in A\cup T-\{\mathds{1}_{G}\}}^{+}$;
  2. (2)
    for all $t\in T$ there exists $h\in H$ s.t. $g=_{G}ht$ and $h$ is non trivial in $H$.
Suppose there exists $i$ and $v\in\Delta_{i}$ as in the hypothesis, then $(t,p)$ with $t\in T$ and $p\in V(\Gamma)$ is such that $(t,p)\cdot g=(t^{\prime},p^{\prime})$ and in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\text{\tiny{$g\mid h$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}$. If $t\neq t^{\prime}$, then in $\mathcal{A}$ we have $t\smash{\stackrel{{\scriptstyle\text{\tiny{$g\mid h$}}}}{{\relbar\joinrel\longrightarrow}}}t^{\prime}\neq t$ hence we have $(t,p)\cdot g=(t^{\prime},p\cdot h)\neq(t,p)$ in $\mathcal{A}\rtimes(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$. Otherwise by definition we have $p\smash{\stackrel{{\scriptstyle\text{\tiny{$h$}}}}{{\relbar\joinrel\longrightarrow}}}p\cdot h\neq p$ in $\Gamma$ and $h$ is non trivial in $H$. So, by Lemma 4.5, there exists $\widetilde{p}\smash{\stackrel{{\scriptstyle\text{\tiny{$\mathring{h}^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(\widetilde{p},x_{0})$ with $\mathring{h}\in S_{A}(h)$. Therefore there exists $p^{\prime}\in V(\Gamma,v)$ for some $i$ and $v\in\Delta_{i}$ such that $p^{\prime}\cdot h\neq p^{\prime}$ and $p^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$h$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\cdot h\subseteq\Gamma(v,x_{0})$. Finally, in $\mathcal{A}\rtimes(\Gamma,v)$ we have $(\mathds{1}_{G},v)\cdot u=(t,p^{\prime})$ where $u$ is such that there exists $\mathds{1}_{G}\smash{\stackrel{{\scriptstyle\text{\tiny{$u\mid z$}}}}{{\relbar\joinrel\longrightarrow}}}t$ in $\mathcal{A}$ with $v\smash{\stackrel{{\scriptstyle\text{\tiny{$z$}}}}{{\relbar\joinrel\longrightarrow}}}p^{\prime}\subseteq(\Gamma,v)$ and $(t,p^{\prime})\cdot g=(t,p^{\prime}\cdot h)\neq(t,p^{\prime})$ since by hyphotesis $tg=_{G}ht$ and $p^{\prime}\cdot h\neq p^{\prime}$. ∎
Remark 4.19.
In particular, this means that any finitely generated subgroup of a virtual direct product of free groups is CF-TR.
4.1. An example
As a concrete example of a transition group of an inverse graph we consider the inverse $A$-digraph $\Omega$ on the alphabet $A=\{a,a^{-1},b,b^{-1},c,c^{-1}\}$ (see Fig. 1). The set of vertices $V(\Omega)=\{p_{i},q_{i},i\in\mathbb{Z}\}$ and the set $E(\Omega)$ of edges is given by:
$p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i+1},\;p_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i+1},\;q_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\quad\forall i\in\mathbb{Z}$
$p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\;p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},\,q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},\;q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},p_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}q_{i},q_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}p_{i},x\in\{c,c^{-1}\}\quad\forall i\in\mathbb{Z}$
We sometimes identify the generic vertex $p_{n}$ with the pair $(n,0)$, $n\in\mathbb{Z}$, while $q_{m}$ with the pair $(m,1)$, $m\in\mathbb{Z}$.
Figure 1. The inverse graph $\Omega$. The loops in red are labelled by $a,a^{-1}$, in blue by $b,b^{-1}$.
Note that $\Omega$ is clearly quasi-transitive and it is quasi-isometric to a tree, thus by [24, Theorem 2] it is a context-free graph. The next proposition characterizes the transition group defined by $\Omega$.
Proposition 4.20.
The transition group of $\Omega$ has the following properties:
  • the subgroup $\langle a,b\rangle$ generated by $a,b$ of the transition group $\mathcal{G}(\Omega)$ is isomorphic to $\mathbb{Z}^{2}$;
  • $\mathcal{G}(\Omega)$ is isomorphic to $G=\langle a,b,c\mid c^{2}=1,ab=ba,ca=bc\rangle$, so it is a semi-direct product $\mathbb{Z}^{2}\rtimes C_{2}$;
  • the action on $V(\Omega)$ is given by the formula $(x,y)\cdot a^{n}b^{m}c^{j}=(x+ym+(1-y)n,\,y+j\mod 2)$ where $a^{n}b^{m}c^{j}$ with $m,n\in\mathbb{Z}$, $j\in\{0,1\}$ is the standard normal form of the semi-direct product.
Proof.
Note that for any vertex $v$ in $\Omega$ we have $v\cdot ab=v\cdot ba$, hence $a,b$ commute in $\mathcal{G}(\Omega)$ ($ab=ba$ in $\mathcal{G}(\Omega)$). Similarly, one can check that for any vertex $v$ we have $v\cdot c^{2}=v$ and $v\cdot(ca)=v\cdot(bc)$. Hence, in $\mathcal{G}(\Omega)$ we also have the equalities $c^{2}=1,ca=bc$ and also $ac=cb$ as a consequence of the involution $c^{2}=1$. Using these equalities, it is easy to see that any element $g\in\mathcal{G}(\Omega)$ is equivalent to a word of the form $g=a^{n}b^{m}c^{j}$ with $m,n\in\mathbb{Z}$, $j\in\{0,1\}$. Now, every element $g\in G$ has a normal form $g=a^{n}b^{m}c^{j}$ with $m,n\in\mathbb{Z}$, $j\in\{0,1\}$. Thus, there is a well defined epimomorphism $\phi:G\to\mathcal{G}(\Omega)$ sending $a^{n}b^{m}c^{j}$ into the corresponding element in $\mathcal{G}(\Omega)$. The formula of the action can be obtained by observing that for a generic vertex $(x,y)\in V(\Omega)$ we have
$(x,y)\cdot a^{n}=(x+(1-y)n,y),\,(x,y)\cdot b^{m}=(x+ym,y),\,(x,y)\cdot c^{j}=(x,\,y+j\mod 2),\,$
Let us show that $\phi$ is injective by using this action. Consider two elements $a^{n}b^{m}c^{j},a^{n^{\prime}}b^{m^{\prime}}c^{j^{\prime}}$, then for any vertex $(x,y)\in V(\Omega)$ the equation $(x,y)\cdot a^{n}b^{m}c^{j}=(x,y)\cdot a^{n^{\prime}}b^{m^{\prime}}c^{j^{\prime}}$ implies $j=j^{\prime}$ and if we choose a vertex with $y=0$ then we conclude that $n=n^{\prime}$, while if we choose $y=1$ we get $m=m^{\prime}$. ∎
5. Detecting torsion
We give some structural results focusing on the existence of torsion elements and on bounds on the order of such elements. Henceforth, we consider a context-free $A$-digraph $\Gamma$ with $N$ end-cone types $\Gamma_{1},\ldots,\Gamma_{N}$ with the corresponding frontier vertices $\Delta_{1},\ldots,\Delta_{N}$ with $\Delta_{i}\subseteq\Gamma_{i}$. We say that two end-cones $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ are self-similar if $\Gamma(p,x_{0})\subseteq\Gamma(q,x_{0})$ and there is an end-isomorphism $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ with $\psi(q)=p$. We have the following lemma.
Lemma 5.1.
Let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic self-similar end-cones of $\Gamma$. Then, for any walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})$ there is an infinite walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\subseteq\Gamma(q,x_{0})$ that is not a circuit.
Proof.
Indeed, let $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ be the end-isomorphism with $\psi(q)=p$. Consider a geodesic $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$ connecting $q$ with $\psi(q)=p$. We claim that also
$q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$
is a geodesic. Any geodesic $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(q,x_{0})$ has to cross a vertex $v\in\Delta(\psi(q),x_{0})$, thus the previous geodesic factors as $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ with $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}v\subseteq\Gamma(q,x_{0})$ and $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(\psi(q),x_{0})$. Since both $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)$ and $\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ are geodesics, respectively in $\Gamma(q,x_{0})$ and $\Gamma(\psi(q),x_{0})$ we deduce $|u_{1}|=|g|$ and $|u_{2}|=|g|$. Hence, $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ is also a geodesic. The same argument can be extended to show that $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{k}(q)\subseteq\Gamma(q,x_{0})$ is a geodesic for any $k\geq 1$. Now, by applying $\psi^{k}$ to the walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$, we deduce that $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\subseteq\Gamma(q,x_{0})$ is an infinite walk that is not a circuit. ∎
To state the next lemma we need to introduce some definitions. For any end-cone type $\Gamma_{i}$ that is end-isomorphic to some end-cone $\Gamma(v,x_{0})$ with $\|v\|_{x_{0}}=n$ there are $k_{i}\leq 2|A|$ second level end-cones $\Gamma(q_{1},x_{0}),\ldots,\Gamma(q_{k_{i}},x_{0})$ contained in $\Gamma(v,x_{0})$ with $\|q_{j}\|_{x_{0}}=n+1$. Roughly speaking we are taking the end-cones of $\Gamma_{i}$ whose frontier vertices are at distance one from $\Delta_{i}$. Put $\gamma=\max\{k_{i},i\in[1,N]\}$ and $\delta=\max\{|\Delta_{i}|,i\in[1,N]\}$. A subgraph $\Lambda$ of $\Gamma$ is called $g$-periodic if $\Lambda$ is equal to the subgraph generated by the circuit $\pi=y\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}y$ for some integer $k\geq 1$. The order $o(g,\Lambda)$ is the smallest integer $m$ such that $\Lambda$ is equal to the subgraph generated by the vertices and edges of the circuit $\pi=y\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{m}$}}}}{{\relbar\joinrel\longrightarrow}}}y$. With a slight abuse of notation, from now on, we identify the subgraph $\Lambda$ with this minimal circuit $\pi$. We have the following lemma.
Lemma 5.2.
For any $g$-periodic subgraph $\pi$ of a context-free graph $\Gamma$ we have $o(g,\pi)\leq|g|\delta\gamma^{N|g|}$.
Proof.
Note that if we cyclically shift $g$ into a word $g^{\prime}$, then $\pi$ is also a $g^{\prime}$-periodic subgraph with $o(g^{\prime},\pi)=o(g,\pi)$. Thus, by a suitable cyclic shift of $g$, we may assume without loss of generality that the subgraph
$\pi=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}y_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots y_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$
with $k=o(g,\pi)$, is contained in the end-cone $\Gamma(y_{1},x_{0})$. Suppose by contradiction that $o(g,\pi)>|g|\delta\gamma^{N|g|}$. We first observe that at most $|g|$ vertices of $\pi$ may be equal to the same vertex $v\in\Gamma$. This is true for $y_{1}$, since $\pi$ is defined to be the minimal circuit. To see this for a generic vertex of $\pi$, we say that a vertex $y_{j}\in\pi$, is of type $g^{\prime}$, where $g^{\prime}$ is a proper prefix of $g$, if there is a sub-walk
$y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{t}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{t}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{j}\subseteq\pi$
for some integer $t\geq 0$. Now, if there are more than $|g|$ vertices of $\pi$ that are equal to the same vertex $v\neq y_{1}$, there would be two vertices, say $y_{i},y_{j}$, $j>i$ such that
$\pi=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{s}g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{g^{\prime\prime}g^{\ell}g^{\prime}}}y_{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime\prime}g^{t}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$
with $g^{\prime}g^{\prime\prime}=g$ and $y_{i}=y_{j}=v$. However, the shorter circuit $y_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{g^{\prime\prime}g^{\ell}g^{\prime}}}y_{j}$ contradicts the minimality of $k$ since $y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\ell+1}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$ would be a circuit with $\ell+1<k$. Let us now estimate the parameter
$\lambda=\max\{\|y_{j}\|_{x_{0}}-\|y_{1}\|_{x_{0}}:y_{j}\in\pi\}$
By the previous remark each vertex in $\pi$ passes from a vertex of $\Gamma$ at most $|g|$ times. Moreover, the possible number of distinct vertices through which the circuit $\pi$ passes is at most $\delta\gamma^{\lambda}$ since there are at most $\delta$ vertices in $\Delta(v,x_{0})$ and each time we are in a vertex in a frontier set, we may choose to move to the second level end-cones of which there are at most $\gamma$. In conclusion, we have that the number of vertices $k$ listed in $\pi$ are at most
$k\leq|g|\delta\gamma^{\lambda}$
which together with the inequality $k>|g|\delta\gamma^{N|g|}$ yields to $\lambda>N|g|$. Now, by the fact that $\Gamma(y_{1},x_{0})$ is quasi isometric to a tree, we have that any sub-walk of $\pi$ starting at some $y_{j}\in\Delta(y_{j},x_{0})$ and ending at $y_{1}$ must pass through some vertex $y_{s}\in\Delta(y_{j},x_{0})$ that belongs to the same set of frontier vertices of $y_{j}$. A noose of $\pi$ at $y_{j}$ is the sub-walk $\Pi(y_{j})$ of $\pi$ starting at $y_{j}$ and ending at $y_{s}\in\Delta(y_{j},x_{0})$ with $s$ minimum. It is clear that $\Pi(y_{j})\subseteq\Gamma(y_{j},x_{0})$. Fix a vertex $\overline{y}$ of $\pi$ with $\|\overline{y}\|_{x_{0}}-\|y_{1}\|_{x_{0}}=\lambda$. We say that a noose $\Pi(y_{j})$ is marked if it contains $\overline{y}$. It is clear that there is a vertex $z_{1}\in\Delta(y_{1},x_{0})$ such that $\Pi(z_{1})$ is a marked noose. There is a marked noose $\Pi(z_{2})$ with $\Pi(z_{2})\subseteq\Pi(z_{1})$ and $\|z_{2}\|_{x_{0}}=\|z_{1}\|_{x_{0}}+1$. Indeed, every sub-walk ending at $\overline{y}$ starting at $z_{1}$ must pass from the frontier vertices $z_{2}$ of a second level end-cone, and if every noose contained in $\Pi(z_{1})$ were not be marked, then $\overline{y}$ would not be contained in $\Pi(z_{1})$, a contradiction. Continuing in this way, we may find a sequence of marked nooses:
$\Pi(z_{1})\supseteq\Pi(z_{2})\supseteq\ldots\supseteq\Pi(z_{\lambda-1})\supseteq\Pi(z_{\lambda})$
where the last one may be an empty walk unless $\Pi(z_{\lambda})$ is of the form $z_{\lambda}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z^{\prime}$ for some $z^{\prime}\in\Delta(z_{\lambda},x_{0})$. Now, let us color each $\Pi(z_{i})$ with the pair $(\Gamma_{s},g^{\prime})$ whenever $z_{i}$ is of type the proper prefix $g^{\prime}$ of $g$ and $\Gamma_{s}$ is the end-cone type with $\Gamma_{s}\simeq_{end}\Gamma(z_{i},x_{0})$. Now, since we have shown before that $\lambda>N|g|$ which is the number of possible colors, we conclude that there are two marked nooses, say $\Pi(z_{i}),\Pi(z_{j})$, $i<j$, colored with the same color, say $(\Gamma_{t},g^{\prime})$. Since $\Pi(z_{j})\subseteq\Pi(z_{i})$, there is a sub-walk
$z_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime\prime}g^{h}g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}z_{j}\subseteq\Pi(z_{i})\subseteq\Gamma(z_{i},x_{0})$
with $g^{\prime}g^{\prime\prime}=g$. Therefore, since $\Gamma(z_{i},x_{0}),\Gamma(z_{j},x_{0})$ are two self-similar end-cones of $\Gamma$, by Lemma 5.1 there is an infinite walk
$z_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{(g^{\prime\prime}g^{h}g^{\prime})^{\omega}}}$
with $z_{i}\in\pi$ that is not a circuit, hence $\pi$ is also $g^{\prime\prime}g^{\prime}$-periodic and $\pi=z_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{(g^{\prime\prime}g^{\prime})^{\omega}}}$, a contradiction. Thus, the bound $o(g,\pi)\leq|g|\delta\gamma^{N|g|}$ holds. ∎
The following proposition gives some information regarding torsion elements.
Proposition 5.3.
The following facts hold:
  • let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic self-similar end-cones of $\Gamma$, and let $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})\subseteq$ be a walk. Then, the element $u\in\mathcal{G}(\Gamma)$ has infinite order.
  • Let $g\in\mathcal{G}(\Gamma)$ be an element of finite order, then the order $o(g)$ of the element $g$ satisfies $\frac{o(g)}{\log_{2}(o(k))}<|g|\delta\gamma^{N|g|}$.
Proof.
The first fact follows from Lemma 5.1 since $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}$ is an infinite walk that is not a circuit we have $u^{i}\cdot q\neq u^{j}\cdot q$ for any $i,j$ with $i\neq j$. Let us prove the bound on the torsion element $g$. Let $o(g)=p_{1}^{\alpha_{1}}\ldots p_{s}^{\alpha_{s}}$ be the prime factorization of $o(g)$. We prove that for any $p_{i}^{\alpha_{i}}$ appearing in the factorization of $o(g)$ we have $p_{i}^{\alpha_{i}}\leq|g|\delta\gamma^{N|g|}$. By definition of the group $\mathcal{G}(\Gamma)$, there is a $g$-periodic subgraph $\pi$ with $o(g,\pi)=m$ and such that $m$ is a multiple of $p_{i}^{\alpha_{i}}$. Thus, by Lemma 5.2 we have
$p_{i}^{\alpha_{i}}\leq m\leq o(g,\pi)\leq|g|\delta\gamma^{N|g|}$
The number of primes appearing in the prime factorization on a positive integer $n$ is maximized by the integers of the form $n=2\cdot 3\cdot\ldots\cdot p_{k}$, thus by a rough estimation $n$ has at most $\log_{2}(n)$ primes appearing in its prime factorization. Hence, the upper bound of the statement follows. ∎
Remark 5.4.
Note that in the case of a finite union of context-free graphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ the first fact of the proposition remains unchanged: whenever there are two self-similar end-cones in one of the graphs and a walk as in the above, the element has infinite order. While for the second fact, instead of considering $\delta$ and $N$, we need $\max_{i\in\{1,\ldots,k\}}\delta_{i}$ and $\max_{i\in\{1,\ldots,k\}}N_{i}$ where $\delta_{i}$ and $N_{i}$ are the corresponding parameters for $\Gamma^{(i)}$.
If the inverse graph $\Gamma$ is infinite, then, since there finitely many end-cone types and by the pigeonhole principle, we may deduce that there are at least two self-similar end-cones. This observation and Proposition 5.3 lead to the following fact.
Proposition 5.5.
The transition group $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})\in{\bf CF-TR}$ of a family of inverse graphs $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ is infinite if and only if at least one graph among $\Gamma^{(1)},\ldots,\Gamma^{(k)}$ is infinite.
Note that given an inverse graph $\Gamma$ presented as an inverse PDA, it is undecidable to check whether $\mathcal{G}(\Gamma)$ is infinite or not. This because it is well-known that it is undecidable to tell whether a PDA is a FSA. Proposition 5.3 gives an alternative proof of the decidability of the torsion problem for the transition groups of context-free graphs. This result may also be derived from the same result for the class of co-context-free groups (see [17, Theorem 9]).
We have the following corollary.
Corollary 5.6.
An infinite group $G=\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})\in{\bf CF-TR}$ does not contain an infinite finitely generated torsion subgroup.
Proof.
This is a consequence of Proposition 4.14, Proposition 5.3 and Proposition 5.5, because any infinite finitely generated subgroup is still in ${\bf CF-TR}$ and so one of the connected components $\Gamma^{(i)}$ of the graph is infinite and it must have an element of infinite order. ∎
Remark 5.7.
By the previous corollary we have that no infinite torsion group is CF-TR, in particular, Grigorchuk group does not belong to such class. However, it is still open the case whether it is CoCF ([12]).
Theorem 5.8.
With the above notation, for a given element $g\in\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$ there is an algorithm to check whether $o(g)$ is finite or not. In other words, the torsion problem is solvable for CF-TR groups.
Proof.
We deal with the case $\mathcal{G}(\Gamma)$, the general case is analog. By Proposition 5.3 if $o(g)$ is finite, then
(2) $\frac{o(g)}{\log_{2}(o(k))}<|g|\delta\gamma^{N|g|}$
Note that $\frac{o(g)}{\log_{2}(o(k))}$, as a function of $o(g)$, has strictly positive derivative for $o(g)>e$, thus for $o(g)\geq 3$ this function is monotonically increasing so there is a (computable) function $C(|g|):\mathbb{N}\to\mathbb{N}$ such that for any $3\leq o(g)<C(|g|)$, the bound (2) is satisfied. Therefore, a generic element $g$ has either finite order satisfying the bound $o(g)<C(|g|)$, or it has infinite order. Hence, to check whether $g$ has finite order it is enough to check if $g^{C(|g|)!}=1$ in $\mathcal{G}(\Gamma)$, or by simply checking whether $g^{k}=1$ for some $k\leq C(|g|)$. Since the word problem for ${\bf CoCF}$ is decidable (see [17, Proposition 7]), the former problem is also decidable. ∎
We know that for $F$ a free group of rank at least $2$, the group $F\times F$ is in CF-TR. Indeed, $F$ is in CF-TR simply by considering the Cayley graph on free generators and, by Proposition 4.15, the class is closed under direct products. This provides an example of a CF-TR group with unsolvable conjugacy problem (see e.g. [21, Theorem 9, Ch.3]) and unsolvable generalized word problem ([20, Theorem 4.6]).
Most of these groups have exponential growth. This happens if the tree which is quasi-isometric to $\Gamma$ has more than two ends. For context-free graphs, this is equivalent to having at least two self-similar end-cones $\Gamma(p,x_{0})\subseteq\Gamma(q,x_{0})$ such that $\Gamma(q,x_{0})$ has $k>1$ second level end-cones.
Proposition 5.9.
If $\Gamma$ is quasi-isometric to a tree with more than two ends, then $\mathcal{G}(\Gamma)$ has exponential growth.
Proof.
Let $V(D_{n}(x_{0}))=\{v\in V(\Gamma):\|v\|_{x_{0}}=n\}$ be the set of vertices $v$ which are at distance $n$ from the root $x_{0}$, and for each $v\in V(D_{n}(x_{0}))$ let us choose a word $w(v)\in A^{*}$ such that $x_{0}\smash{\stackrel{{\scriptstyle\text{\tiny{$w(v)$}}}}{{\relbar\joinrel\longrightarrow}}}v$ is a walk connecting $x_{0}$ with $v$. Let
$G_{n}=\{g_{v}\ \in\mathcal{G}(\Gamma):g_{v}\mbox{ is represented by }w(v)\mbox{ for each }v\in V(D_{n}(x_{0}))\}$
be the set of elements of $\mathcal{G}(\Gamma)$ corresponding to $w(v)$ for each $v\in V(D_{n}(x_{0}))$. Note that $|G_{n}|=|V(D_{n}(x_{0}))|$ since if $v_{1},v_{2}\in V(D_{n}(x_{0}))$ are two distinct vertices, then the corresponding elements $g_{v_{1}},g_{v_{2}}$ in $\mathcal{G}(\Gamma)$ are different since they acts differently on $x_{0}$. Moreover, each $g_{v}\in G_{n}$ is in the ball $B_{n}$ of radius $n$ in the Cayley graph of $\mathcal{G}(\Gamma)$. Thus, for each $n>0$ we have the exponential upper bound
$B_{n}\geq|G_{n}|=|V(D_{n}(x_{0}))|$
Remark 5.10.
If we are dealing with a finite union of graphs, we ask that at least one of them is quasi-isometric to a tree with more than two ends.
6. Poly-context-free groups and co-context-free transition groups
In this section, we try to shed some light on the first question of Problem 4.11. At the end of the section we also make some considerations on some conjectures regarding poly-context-free groups.
Throughout this section, we focus on groups whose word problem is the intersection of finitely many inverse-context-free languages, subsequently denoted as Poly-ICF. The following theorem shows that this class is contained in CF-TR.
Theorem 6.1.
Let $G$ be a group whose word problem $WP(G;A)$ is the intersection of $k$ context-free languages accepted by rooted inverse graphs:
$WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma^{(i)},x_{i})$
Then, $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is in CF-TR.
Proof.
The $A$-digraph
[…]
with $\alpha=(x_{1},\ldots,x_{k})$ is clearly inverse since each $\Gamma_{i}$ is inverse, and since $WP(G;A)=L(\Gamma,\alpha)$, by Proposition 2.5 we deduce that $\Gamma$ is isomorphic to the Cayley graph $\operatorname{Cay}(G;A)$. We recall that for any $u\in A^{*}$ and $x\in V(\Gamma)$, $x\cdot u$ is the endpoint of the walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}(x\cdot u)$ in an inverse and complete $A$-digraph $\Gamma$. We clearly have $\alpha\cdot u=(x_{1}\cdot u,\ldots,x_{k}\cdot u)$, furthermore, it is straightforward to check that the following equality holds for all $u\in A^{*}$:
[…]
Now, since $\Gamma$ is a Cayley graph $(\Gamma,\alpha)\simeq(\Gamma,\alpha\cdot u)$, thus for all $u\in A^{*}$ we have $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma,x_{i}\cdot u)$. Hence, by recalling equation (1), we conclude that
(3) $WP(G;A)=\bigcap_{i=1}^{k}\left(\bigcap_{y\in V(\Gamma^{(i)})}L(\Gamma^{(i)},y)\right)=\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})=\mathcal{L}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$
i.e., $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$. ∎
An interesting question is whether there exists examples of CF-TR groups which are not poly-context-free. The next result tells us that such examples needs to be non-quasi-transitive (for a specific example, refer to Section 10). More precisely, it gives a characterization that offers a kind of converse to the previous theorem.
Theorem 6.2.
Let $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$ be a collection of quasi-transitive context-free inverse $A$-digraphs. Then, $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is a group in CF-TR if and only if $G$ is virtually a finitely generated subgroup of the direct product of free groups. In particular, it belongs to the class of poly-context-free groups whose word problem is the intersection of finitely many languages accepted by rooted quasi-transitive context-free inverse graphs.
Proof.
Fix a generic inverse $A$-digraph $\Gamma^{(i)}$ of the family $\{\Gamma^{(1)},\ldots,\Gamma^{(k)}\}$. Since it is quasi-transitive, there is a finite complete set $F_{i}\subseteq V(\Gamma^{(i)})$ of representatives of the equivalence relation given by the orbits of $\operatorname{Aut}(\Gamma_{i})$. Since by Proposition 2.5 we have that for any $\varphi\in\operatorname{Aut}(\Gamma^{(i)})$, $L(\Gamma^{(i)},v)=L(\Gamma^{(i)},\varphi(v))$ we deduce:
$\mathcal{L}(\Gamma^{(i)})=\bigcap_{p\in V(\Gamma^{(i)})}L(\Gamma^{(i)},p)=\bigcap_{\varphi\in\operatorname{Aut}(\Gamma^{(i)}),p\in F_{i}}L(\Gamma^{(i)},\varphi(p))=\bigcap_{p\in F_{i}}L(\Gamma^{(i)},p)$
Since the word problem $WP(G;A)$ for $G=\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is the intersection $\bigcap_{i=1}^{k}\mathcal{L}(\Gamma^{(i)})$ we have that $\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is a poly-context-free group. In particular
$WP(G;A)=\bigcap_{i=1}^{k}\left(\bigcap_{p\in F_{i}}L(\Gamma^{(i)},p)\right)$
is the intersection of finitely many context-free languages accepted by quasi-transitive inverse $A$-digraphs, hence by [24, Theorem 3] this is equivalent for the group $G$ to be virtually a finitely generated subgroup of the direct product of free groups. Conversely, by [24, Proposition 12] if $G$ is virtually a finitely generated subgroup of the direct product of free groups, then the word problem $WP(G;A)=\bigcap_{i=1}^{k}L(\Gamma^{(i)},x_{i})$ is the intersection of finitely many inverse-context-free languages accepted by inverse graphs that are quasi-transitive. Hence, by Theorem 6.1 we have that the group $G\simeq\mathcal{G}(\Gamma^{(1)}\sqcup\ldots\sqcup\Gamma^{(k)})$ is the transition group of the disjoint union of finitely many inverse context-free $A$-digraphs that are quasi-transitive. ∎
The previous theorem offers another characterization of groups that are virtually finitely generated subgroups of the direct product of free groups. Indeed, in [24, Theorem 3.1] it is shown that such class is also characterized by the groups whose word problem is in Poly-QT-ICF, i.e., languages that are the intersection of finitely many languages accepted by inverse graphs that are quasi-transitive and quasi-isometric to a tree. Note that the example presented in subsection 4.1 is in Poly-QT-ICF since by a quick inspection of Fig. 1 shows that $\Omega$ is quasi-transitive.
Following [8] we give the following construction. For $c=\langle c_{0},\ldots,c_{s}\rangle\in\mathbb{Z}^{s+1}$ with $s\geq 1,c_{0},c_{s}\neq 0$ and $\mathrm{GCD}(c_{0},\ldots,c_{s})=1$, the group $G(c)$ is defined by the presentation $\langle a,b\mid\mathfrak{R}_{c}\rangle$, where $\mathfrak{R}_{c}=\{[b,b^{a^{i}}](i\in\mathbb{Z}),b^{c_{0}}(b^{a})^{c_{1}}\ldots(b^{a^{s}})^{c_{s}}\}$ via the notation $x^{y}:=y^{-1}xy$ and $[x,y]:=x^{-1}y^{-1}xy$. These groups are called $Gc$-groups and we say that a $Gc$-group is proper if it is not virtually abelian.
We end this section with some consideration on Brough Conjecture 2 where it is conjectured that a finitely generated poly-context-free group is solvable if and only if it is virtually abelian. In this direction we have the following proposition.
Proposition 6.3.
If $G\in{\bf CF-TR}$ is a finitely generated transition group that is solvable, then one of the following holds:
  • $G$ is virtually abelian;
  • $G$ contains the free abelian group $\mathbb{Z}^{\infty}$ of countably infinite rank.
  • $G$ has arbitrarily large finite subgroups.
Proof.
Combining results from [8], Theorem 2.12 in [9] shows that a solvable finitely generated group $G$ which is not virtually abelian either contains
  1. (1)
    $\mathbb{Z}^{\infty}$, or
  2. (2)
    a proper $Gc$-group, or
  3. (3)
    a finitely generated subgroup $H$ with an infinite normal torsion subgroup $U$, such that $H/U$ is either free abelian or a proper $Gc$-group.
By assumption the group $G$ is CF-TR then it is CoCF and so, since by [9, Proposition 4.13] a $Gc$-group is poly-context-free or co-context-free if and only if it is virtually abelian, we deduce that case (2) is not possible. Let us now analyze case (3) and assume that $G$ contains a subgroup $H$ with an infinite normal torsion subgroup $U$. Suppose that $U=\langle g_{i}:i\in\mathbb{N}\rangle$ and consider the subgroups $U_{n}=\langle g_{i}:i\leq n\rangle\leq U$. The subgroups $U_{n}$ are clearly finitely generated torsion subgroups of $G$ and thus, by Corollary 5.6, they are all finite. Clearly, the order of the subgroups $U_{n}$ becomes arbitrarily large since $U$ is infinite. ∎
In Brough Conjecture 3 it is also conjectured that a poly-context-free group does not have arbitrarily large finite subgroups. In the intersection of CF-TR with poly-context-free groups we have the following fact.
Corollary 6.4.
If it is true that a group $G$ which is both poly-context-free and CF-TR does not have arbitrarily large finite subgroups, then the following are equivalent for a finitely generated group
  1. (1)
    $G$ is virtually abelian
  2. (2)
    $G$ is a solvable poly-context-free transition group in ${\bf CF-TR}$
Proof.
(1) $\implies$ (2) Since $\mathbb{Z}/n\mathbb{Z}$ and $\mathbb{Z}$ can be realized as transition groups of context-free graphs by Example 4.2, then Proposition 4.15 and Proposition 4.18 imply that finitely generated virtually abelian groups are also in the class of transition groups as well as poly-context-free by [9]. (2) $\implies$ (1) The converse implication follows by Proposition 6.3 and the fact that $\mathbb{Z}^{\infty}$ is not poly-context-free, as it contains free abelian subgroups of rank $k$ for all $k\in\mathbb{N}$, see [9, Lemma 4.6]. ∎
From the previous discussion we have the following natural questions.
Question 6.5.
ciao
  • Is it possible to realize a solvable transition group associated to a context-free graph containing $\mathbb{Z}^{\infty}$ as a subgroup? In particular, is the group described in Definition 10.4 solvable?
  • Is it true that a transition group associated to a context-free graph does not have arbitrarily large finite subgroups? By the argument at the end of the proof of Proposition 6.3 this is equivalent to not having an infinite torsion subgroup.
7. Perturbing a graph: a structural result
A natural question about these groups is how much small local changes in the graph affect the associated transition groups. In this section we investigate what happens to the transition group when we locally perturb a context-free graph $(\Gamma,x_{0})$. The idea is to consider a ball $D_{n}(x_{0})$ and to associate with the inverse graph $\Lambda_{n}=\Gamma\setminus D_{n}(x_{0})$ a group called the boundary group $\mathrsfs{B}_{n}$ at level $n$, and then to give a structural result connecting this group to the transition group $G=\mathcal{G}(\Gamma)$.
Definition 7.1 (Boundary group).
Consider the following subset of $A^{*}$
$\mathcal{O}_{n}=\{u\in A^{*}:\mbox{ if }x\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x^{\prime}\mbox{ is a walk in }\Lambda_{n}\mbox{ then }x=x^{\prime}\}$
of the words that either label a circuit totally contained in $\Lambda_{n}$, or whose walk intersects $S_{n}(x_{0})$. It is easy to verify that $\mathcal{O}_{n}$ is a submonoid of $A^{*}$ satisfying the property that for any $u\in A^{*}$ and $w\in\mathcal{O}_{n}$ we have $u^{-1}wu\in\mathcal{O}_{n}$. Therefore, if $\pi:A^{*}\to G$ is the natural map, the image $\mathrsfs{B}_{n}=\pi(\mathcal{O}_{n})$ is a normal subgroup of $G$ called the boundary group at level $n$.
We have the following key lemma.
Lemma 7.2.
With the above notation, for any $u\in\mathcal{O}_{n}$ and $x\in V(\Gamma)$ we have $x\cdot u^{k}=x$ for some $k\leq\left|S_{n}(x_{0})\right|\,\left|u\right|$.
Proof.
Consider the infinite walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$, and let us denote by $x_{i}^{(j)}$ the vertices of such walk with $x=x_{0}^{(0)}$ and $x_{i}^{(j)}$ for $0\leq i\leq|u|$ are all the vertices between $x_{0}^{(j)}$ and $x_{0}^{(j+1)}$ so that $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ are the corresponding subwalks of $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$. If $x\cdot u=x$ we are done, thus we may suppose $x\cdot u\neq x$. Note that every subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ intersects $S_{n}(x_{0})$. Indeed, if $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ were totally contained in $\Lambda_{n}$, then since $u\in\mathcal{O}_{n}$, we would have $x_{0}^{(j)}=x_{0}^{(j+1)}$, hence $x\cdot u=x$ when $j=0$, a contradiction. Now if we have that $x_{i}^{{(j)}}=x_{i}^{{(s)}}=v\in S_{n}(x_{0})$, for some $s\neq j$, then by the determinism of $\Gamma$ we have $x\cdot u^{k}=x$ for some $k$. Since the walk $x\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ is infinite and $S_{n}(x_{0})$ is a finite set. Either one of the two cases happens. To give the upper bound on $k$, we consider worst case scenario: each subwalk $x_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}x_{0}^{(j+1)}$ contains at most a vertex of $S_{n}(x_{0})$, therefore by the pigeonhole principle it is easy to see that condition $x_{i}^{{(j)}}=x_{i}^{{(s)}}$ occurs for some $j,s\leq\left|S_{n}(x_{0})\right|\,\left|u\right|$. ∎
As an immediate consequence of the previous lemma, we have the following corollary.
Corollary 7.3.
For all $n\geq 1$, the boundary group $\mathrsfs{B}_{n}$ at level $n$ is a torsion normal subgroup of $G$ where each element $g$ has order $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$ (where ||$\cdot$|| ⋅ | is the group metric with respect to $A$).
In view of the structural result contained in Proposition 6.2 an interesting case is when the graph $\Lambda_{n}$ can be obtained by a quasi-transitive inverse graph by cutting out the disk $D_{n}(x_{0})$, more precisely we define the following class of inverse graphs.
Definition 7.4.
An inverse graph $(\Gamma,x_{0})$ is called locally-quasi-transitive if there is a quasi-transitive inverse graph $(\Theta,y_{0})$ and an integer $n\geq 1$ such that $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$.
In case all the orbits are infinite we may explicitly calculate the set $\mathcal{O}_{n}$.
Lemma 7.5.
Suppose the quasi-transitive graph $\Theta$ has finitely many orbits $O_{1},\ldots,O_{m}$ which are all infinite. Let $R=\{p_{1},\ldots,p_{m}\}$ be a set of representative vertices such that $p_{i}\in O_{i}$ for $1\leq i\leq m$. Then,
$\mathcal{O}_{n}=\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$
Proof.
The equality $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ follows by Proposition 2.5 as two vertices in the same orbits have the same associate language. Since $\bigcap_{p\in R}L(\Theta,p)=\mathcal{L}(\Theta)$ are words that label a circuit at every point, we clearly have the inclusion $\bigcap_{p\in R}L(\Theta,p)\subseteq\mathcal{O}_{n}$. Conversely, take a word $u\in\mathcal{O}_{n}$, for each $p_{i}\in R$ we claim that $u$ labels a circuit at $p_{i}$. Indeed, since the orbit $O_{i}$ is infinite, we may find a vertex $p_{i}^{\prime}\in O_{i}$ that is at sufficiently large distance from the disk $D_{n}(x_{0})$ such that $p_{i}^{\prime}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is totally contained in $\Lambda_{n}$, hence by definition $z=p_{i}^{\prime}$ and therefore $u$ labels also a circuit at $p_{i}$. Therefore, $u\in\bigcap_{p\in R}L(\Theta,p)$ showing the other inclusion. ∎
In case the inverse graph $(\Gamma,x_{0})$ is locally-quasi-transitive where the associated quasi-transitive graph $\Theta$ has all the orbits that are infinite, we say that $(\Gamma,x_{0})$ is locally-quasi-transitive with infinite orbits. We are now in position to prove the following structural theorem.
Theorem 7.6.
Let $(\Gamma,x_{0})$ be a context-free inverse graph that is also locally-quasi-transitive with infinite orbits and let $G=\mathcal{G}(\Gamma)$ its transition group. Then, there is a short exact sequence
$\mathds{1}\rightarrow\mathrsfs{B}_{n}\rightarrow G\rightarrow H\rightarrow\mathds{1}$
where $H$ is a group that is virtually a finitely generated subgroup of the direct product of free groups and $\mathrsfs{B}_{n}$ is a torsion normal subgroup where the order of each element $g$ is bounded above by $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$. In particular we have $H\simeq G/_{\mathrsfs{B}_{n}}$.
Proof.
Let $\Theta$ be the inverse graph which is quasi-transitive with infinite orbits, such that there exists an integer $n\geq 1$ with $\Gamma\setminus D_{n}(x_{0})=\Lambda_{n}=\Theta\setminus D_{n}(y_{0})$. Since there are finitely many end-cone types with frontier vertices contained inside the ball of radius $n$, we have that $\Gamma$ and $\Theta$ differ from finitely many end-cones up to end-isomorphism. Hence, $\Theta$ is also context-free. Therefore, by Proposition 6.2 we have that $H=\mathcal{G}(\Theta)$ is virtually a finitely generated subgroup of the direct product of free groups. Now, by Lemma 7.5 we have $\mathcal{O}_{n}=\mathcal{L}(\Theta)$, hence since $\mathcal{L}(\Gamma)\subseteq\mathcal{O}_{n}=\mathcal{L}(\Theta)$, there is a well-defined epimorphism $\psi:\mathcal{G}(\Gamma)\rightarrow\mathcal{G}(\Theta)$, with $\ker(\psi)=\pi(\mathcal{O}_{n})=\mathrsfs{B}_{n}$ ($\pi$ is the natural map from $A^{*}$ to the group $\mathcal{G}(\Gamma)$). From which we have the short exact sequence in the statement and the fact that $H\simeq G/_{\mathrsfs{B}_{n}}$. Moreover, by Corollary 7.3, we have that the boundary group $\mathrsfs{B}_{n}$ is a torsion subgroup of $\mathcal{G}(\Gamma)$ where the order of each element $g$ is bounded above by $o(g)\leq\left|S_{n}(x_{0})\right|\,\left|g\right|$. ∎
It is worth mentioning that in the settings of the previous proposition, the boundary group $\mathrsfs{B}_{n}$ is generated by a poly-context-free language. In Section 10 we will show a concrete example of a transition group associated to a context-free locally-quasi-transitive graph having boundary group which is infinite and whose elements have order two. This gives a negative answer to the second point of Question 6.5.
8. Rationality of CF-TR
An asynchronous transducer is a quintuple $T=(Q,q_{0},\Sigma,\delta,o)$, where
  1. (1)
    $Q$ is a finite set states;
  2. (2)
    $q_{0}\in Q$ is the initial state;
  3. (3)
    $\delta$ is a function $Q\times\Sigma\rightarrow Q$ called transition function;
  4. (4)
    $o$ is a function $Q\times\Sigma\rightarrow\Sigma^{*}$ called output function.
For the sake of brevity, we denote $\delta(q,\sigma)=\widetilde{q}$ and $o(q,\sigma)=w$ simply with $q\smash{\stackrel{{\scriptstyle\text{\tiny{$\sigma|w$}}}}{{\relbar\joinrel\longrightarrow}}}\widetilde{q}$. It is straightforward that one can extend $o$ and $\delta$ to finite and infinite words. More precisely, if $\widehat{\Sigma}=\Sigma^{*}\cup\Sigma^{\omega}$, we have $\widehat{o}:Q\times\widehat{\Sigma}\rightarrow\widehat{\Sigma}$ and $\widehat{\delta}:Q\times\widehat{\Sigma}\rightarrow Q$ simply by the formula
$\widehat{o}(q,\sigma_{1}\sigma_{2})=o(q,\sigma_{1})o(\delta(q,\sigma_{1}),\sigma_{2})\text{ and }\widehat{\delta}(q,\sigma_{1}\sigma_{2})=\delta(\delta(q_{1},\sigma_{1}),\sigma_{2}).$
Furthermore, fixed $q_{0}$, we have the functions $q_{0}\ast:\widehat{\Sigma}\rightarrow\widehat{\Sigma}$ and $q_{0}\cdot:Q\rightarrow Q$ with
$q_{0}\ast w=\widehat{o}(q_{0},w)\text{ and }q_{0}\cdot w=\widehat{\delta}(q_{0},w).$
A homeomorphism $\varphi:\Sigma^{\omega}\rightarrow\Sigma^{\omega}$ is rational if there exists an asynchronous transducer $(Q,q_{0},\Sigma,\delta,o)$ such that $\varphi(\eta)=q_{0}\ast\eta$ for all $\eta\in\Sigma^{\omega}$. In [14], it has been proved that the set of rational homeomorphisms over an alphabet $\Sigma$ is indeed a group $\mathcal{R}_{\Sigma}$ (called the rational group) and any two rational groups over two different alphabets are isomorphic (so it makes sense to simply write $\mathcal{R}$). In [4] it is shown that the rational group $\mathcal{R}$ is simple and non-finitely generated. We say that a group which embeds in any rational group is rational. From a topological point of view $\Sigma^{\omega}$ is a Cantor space where the metric is given by the longest common prefix. In this way, there is a well-defined notion of bi-lipschitz homeomorphism. A transducer associated to a homeomorphism is said to be bi-lipschitz if the homeomorphism is bi-lipschtz. The family of groups which can be represented by bi-lipschitz transducers includes Thompson-like groups, self-similar groups and hyperbolic groups (see [3]). We want to construct a bi-lipschitz transducer for each element of our group belonging to CF-TR, so that any such group embeds in $\mathcal{R}$. To fix the notation, let $\{\Gamma_{0},\ldots,\Gamma_{N}\}$ be the representative of each end-isomorphim class of the context-free inverse graph $(\Gamma,x_{0})$, and let $\{\Delta_{0},\ldots,\Delta_{N}\}$ be the corresponding frontier vertices. We put $\Gamma_{0}=\Gamma$. As we have discussed at the end of Lemma 5.1, for each end-cone $\Gamma(x,x_{0})$ of a context-free graph inverse graph $(\Gamma,x_{0})$ we have a certain number of second-level end-cones which are end-isomorphic to some representative among $\{\Gamma_{0},\ldots,\Gamma_{N}\}$. This means that each $\Gamma_{i}$ has $k(i)$ second-level end-cones that we denote by $\Gamma^{(i)}_{1},\ldots,\Gamma^{(i)}_{k(i)}$. Each subgraph $\Gamma^{(i)}_{j}$ is end isomorphic to some end-cone, say $\Gamma_{h}\in\{\Gamma_{0},\ldots,\Gamma_{n}\}$, and let us denote by $\Psi(i,j):\Gamma^{(i)}_{j}\to\Gamma_{h}$ such end-isomorphism and put $\psi(i,j)=h$, where $i\in[0,N]$, $j\in[1,k(i)]$. We also denote by $\Delta_{j}^{(i)}$ the frontier vertices of $\Gamma^{(i)}_{j}$. We now define the alphabet on which our transducer will act:
$\Lambda=\{(\Gamma_{j}^{(i)},v):v\in\Delta_{j}^{(i)},i\in[1,N],j\in[1,k(i)]\}\cup\{(\Gamma_{j}^{(i)}),i\in[1,N],j\in[1,k(i)]\}\cup\{(\Gamma_{0}),(\Gamma_{0},x_{0})\}$
We say that a letter $\lambda\in\Lambda$ is final if it is of the form $\lambda=(\Gamma_{j}^{(i)},v)$; in this case we put $\lambda(1)=\Gamma_{j}^{(i)}$, $\lambda(2)=v$, while for a non-final letter $\lambda=(\Gamma_{j}^{(i)})$ we put $\lambda(1)=\Gamma_{j}^{(i)}$, $\lambda(2)=0$. Moreover, for any $\Gamma_{j}^{(i)},\Gamma_{k}^{(h)}$ we say that $\Gamma_{j}^{(i)}<\Gamma_{k}^{(h)}$ if $\Gamma_{k}^{(h)}$ is a second-level end-cone of $\Psi(i,j)(\Gamma_{j}^{(i)})$. Note that in this case we have $h=\psi(i,j)$. We extend this relation of the letters of $\Lambda$ by putting $\lambda_{1}<\lambda_{2}$ if $\lambda_{1}(1)<\lambda_{2}(1)$. A word $w\in\Lambda^{*}$ is called well-formed, if $w=\lambda_{0}\ldots\lambda_{\ell}$ with $\lambda_{0}=(\Gamma_{0})$, $\lambda_{\ell}$ is the only final letter of $w$, and $\lambda_{0}<\lambda_{1}<\ldots<\lambda_{\ell}$; we denote by $\mathcal{F}(\Lambda)$ the set of well-formed words on $\Lambda$. Roughly speaking, a well-formed word $w$ represents a vertex $x\in\Gamma$ and the sequence $\lambda_{0}(1)<\lambda_{1}(1)<\ldots<\lambda_{\ell}(1)$ represents the sequence of end-cones traversed by a geodesic connecting $x_{0}$ to $x$. More precisely, an end-geodesic is a sequence of end-cones $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ such that $\|x_{i+1}\|_{x_{0}}=\|x_{i}\|_{x_{0}}+1$ for $i\in[0,\ell-1]$. Note that each $\Gamma(x_{i+1},x_{0})$ is a second-level end-cone of $\Gamma(x_{i},x_{0})$. A rooted end-geodesic is given by end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0}),x_{\ell})$ where we fix a vertex $x_{\ell}$ in the last end-cone. Henceforth, when we write $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ we always mean a rooted end-geodesic with root given by the last vertex $x_{\ell}$ in the sequence. We have the following lemma.
Lemma 8.1.
Let $w=(\Gamma_{0})(\Gamma_{j_{1}}^{(i_{1})})\ldots(\Gamma_{j_{s}}^{(i_{s})})\ldots(\Gamma_{j_{\ell-1}}^{(i_{\ell-1})})(\Gamma_{j_{\ell}}^{(i_{\ell})},v_{\ell})\in\mathcal{F}(\Lambda)$ be a well-formed word in $\Lambda^{*}$. Then, there is a unique rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ and end-isomorphisms
$\varphi_{s}:\Gamma(x_{s},x_{0})\to\Gamma_{j_{s}}^{(i_{s})},\,s\in[1,\ell]$
where $\varphi_{0}$ is the identity map, $\varphi_{\ell}(x_{\ell})=v_{\ell}$, and satisfying $(\Psi(i_{s-1},j_{s-1})\circ\left.\varphi_{s-1})\right|_{\Gamma(x_{s},x_{0})}=\varphi_{s}$. Vice versa, for any rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ there is a unique well-formed word satisfying the previous properties.
Proof.
By induction on the parameter $s$. If $s=0$, then $\varphi_{0}$ the identity satisfies the desired conditions. Since $w$ is well-formed, $\Gamma_{j_{1}}^{(i_{1})}$ is a second-level end-cone that is a subgraph of $\Gamma_{0}$. In this case, necessarily $\Gamma(x_{1},x_{0})=(\Gamma_{j_{1}}^{(i_{1})},x_{1})$ and we put $\varphi_{1}=\left.\varphi_{0}\right|_{\Gamma(x_{1},x_{0})}$. Let us assume that we have proved that there is an rooted end-geodesic
$\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{s-1},x_{0})$
and end-isomorphisms $\varphi_{i}$, $i\in[1,s]$, satisfying the statement of the lemma. Now, consider the end-isomorphism $\Psi(i_{s-1},j_{s-1}):\Gamma_{j_{s-1}}^{i_{s-1}}\to\Gamma_{h}$, $h=\psi(i_{s-1},j_{s-1})$ and consider the end-isomorphism $\phi_{s}=\Psi(i_{s-1},j_{s-1})\circ\varphi_{s-1}:\Gamma(x_{s-1},x_{0})\to\Gamma_{h}$. Since $w$ is well-formed, $\Gamma_{j_{s}}^{(i_{s})}$ is a second-level end-cone of $\Gamma_{h}$. Therefore, there is a second-level end-cone $\Gamma(x_{s},x_{0})=(\phi_{s}^{-1}(\Gamma_{j_{s}}^{(i_{s})}),\phi_{s}^{-1}(v_{s}))$ of $\Gamma(x_{s-1},x_{0})$, for some $v_{s}\in\Delta_{j_{s}}^{(i_{s})}$. Now, by putting $x_{s}=\phi_{s}^{-1}(v_{s})$ and $\varphi_{s}=\left.\phi_{s}\right|_{\Gamma(x_{s},x_{0})}$ we have that $\Gamma(x_{s},x_{0})$ is the unique second-level end-cone of $\Gamma(x_{s-1},x_{0})$ satisfying $\varphi_{s}(x_{s})=v_{s}$ and $(\Psi(i_{s-1},j_{s-1})\circ\left.\varphi_{s-1})\right|_{\Gamma(x_{s},x_{0})}=\varphi_{s}$.
For the other direction, let $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ be a rooted end-geodesic. Let $\varphi_{0}$ be the identity map, and assume that we have constructed a word
$(\Gamma_{0})(\Gamma_{j_{1}}^{(i_{1})})\ldots(\Gamma_{j_{s}}^{(i_{s})})\ldots(\Gamma_{j_{\ell-1}}^{(i_{\ell-1})},v_{\ell-1})$
and end-isomorphisms $\varphi_{i}$, $i\in[0,\ell-1]$ satisfying the statement of the lemma. Then, let $\Gamma_{j_{\ell}}^{(i_{\ell})}$ to be the second-level end-cone $\Psi(i_{\ell-1},j_{\ell-1})\circ\varphi_{\ell-1}(\Gamma(x_{\ell},x_{0}))$ of $\Gamma_{j_{\ell-1}}^{(i_{\ell-1})}$, and $v_{\ell}=\Psi(i_{\ell-1},j_{\ell-1})\circ\varphi_{\ell-1}(x_{\ell})$. Then,
$(\Gamma_{0})(\Gamma_{j_{1}}^{(i_{1})})\ldots(\Gamma_{j_{s}}^{(i_{s})})\ldots(\Gamma_{j_{\ell-1}}^{(i_{\ell-1})})(\Gamma_{j_{\ell}}^{(i_{\ell})},v_{\ell})$
satisfies the statement of the lemma, and the associated rooted end-geodesic is exactly $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$. ∎
By the previous Lemma 8.1 for each well-formed word
$w=(\Gamma_{0})(\Gamma_{j_{1}}^{(i_{1})})\ldots(\Gamma_{j_{s}}^{(i_{s})})\ldots(\Gamma_{j_{\ell-1}}^{(i_{\ell-1})})(\Gamma_{j_{\ell}}^{(i_{\ell})},v_{\ell})\in\mathcal{F}(\Lambda)$
we may associate a vertex $x_{\ell}=\varphi_{\ell}^{-1}(v_{\ell})$, and vice versa, given a vertex $x_{\ell}$ in $\Gamma$, there is a unique rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ for which we may associate a unique well-formed word that we denote by $w[x_{\ell}]$. Therefore, there is a one to one correspondence between vertices of $\Gamma$ and well-formed words. For any word $w\in\Lambda^{*}$, let $w\Lambda^{\omega}$ be the clopen set formed by the infinite words starting with the word $w$. By, Lemma 8.1 we have the following fact.
Lemma 8.2.
The map $\Xi$
$\Xi:V(\Gamma)\to\{w\Lambda^{\omega}:w\in\mathcal{F}(\Lambda)\}$
from the vertices of $\Gamma$ to the set of infinite words starting with a well-formed word defined by $\Xi(x)=w[x]\Lambda^{\omega}$ is a bijection.
This is a key point that will be used in proving that the action of each letter $a\in A$ on $\Gamma$ is rational. We now define the asynchronous transducer $\mathcal{T}_{\Gamma}$. This transducer acts on $\Lambda$ and has set of states
$Q_{\Gamma}=\left\{(\Gamma_{j}^{(i)},a):i\in[1,N],j\in[1,k(i)],a\in A\right\}\cup\left\{(\Gamma_{0},a):a\in A\right\}\cup\{E\}$
We now describe the set of transitions. The state $E$ is a sink state, which means that we have the transitions $E\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda\mid\lambda$}}}}{{\relbar\joinrel\longrightarrow}}}E$ for all $\lambda\in\Lambda$. For each fixed $a\in A$ we now describe a transducer that mimics the action of the letter $a$ when we are in a vertex $x$ encoded in the closed set $\Xi(x)$. This is done by first checking whether the word in the input encodes a well-formed word representing a vertex $x$, and then changing the last letters to match the new vertex $x\cdot a$ in the new encoding in the output. The transitions are defined as follows. We start by describing how the initial state acts:
(4) $\begin{cases}(\Gamma_{0},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|1$}}}}{{\relbar\joinrel\longrightarrow}}}(\Gamma_{0},a)&\mbox{if }\lambda=(\Gamma_{0})\\ (\Gamma_{0},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda$}}}}{{\relbar\joinrel\longrightarrow}}}E&\mbox{ in the other cases when }\lambda\mbox{ is not final. }\end{cases}$
where the case when $\lambda$ is final is treated later in (7) and (8). The transition of a generic state $(\Gamma_{j}^{(i)},a)$ with input $\lambda$ depends on whether $\lambda(1)=\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)}$ and whether $\lambda$ is final or not. Let us first assume $\lambda(1)=\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)}$. Now, if $\lambda$ is not final we have the following transition:
(5) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}(\lambda(1),a),\mbox{ where }\lambda^{\prime}=(\Gamma_{j}^{(i)}),\mbox{\bf if }\lambda(2)=0\mbox{ and }\lambda(1)<\Gamma_{j}^{(i)};\$
in case it is final, we are in three cases (6), (7), (8) described below and the action depends on the end-cone $\lambda(1)=\Gamma_{s}^{(h)}$ and the vertex $\lambda(2)$. In the first case, when applying the letter $a$ to the vertex $\lambda(2)$ we get back to the previous end-cone:
(6) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}E,\lambda^{\prime}=(\Gamma_{j}^{(i)},z),\mbox{ \bf if }\lambda(2)\neq 0,\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)},$
[…]
The second case is when applying the letter $a$ to the vertex $\lambda(2)$ we are staying in the same frontier vertices of the end-cone $\lambda(1)=\Gamma_{s}^{(h)}$:
(7) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}\lambda^{\prime\prime}$}\,}}}{{\relbar\joinrel\relbar\joinrel\longrightarrow}}}E,\mbox{ where }\lambda^{\prime}=(\Gamma_{j}^{(i)}),\lambda^{\prime\prime}=(\Gamma_{s}^{(h)},z),\mbox{ \bf if }\lambda(2)\neq 0,\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)}$
[…]
The last case occurs when applying the letter $a$ to the vertex $\lambda(2)$ we are moving to a frontier vertex on a second-level end-cone $\Gamma_{t}^{(m)}$ of $\lambda(1)=\Gamma_{s}^{(h)}$:
(8) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}\lambda^{\prime\prime}\lambda^{\prime\prime\prime}$}\,}}}{{\relbar\joinrel\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}}}E,\mbox{ where }\lambda^{\prime}=(\Gamma_{j}^{(i)}),\lambda^{\prime\prime}=(\Gamma_{s}^{(h)}),\lambda^{\prime\prime\prime}=(\Gamma_{t}^{(m)},z)\mbox{ \bf if }\lambda(2)\neq 0,$
$\Gamma_{t}^{(m)}<\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)},\mbox{ and }\Psi(s,h)(\lambda(2))\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z\mbox{ is a walk in }\Psi(s,h)(\Gamma_{s}^{(h)})=\Gamma_{m},\mbox{ with }z\in\Delta_{t}^{(m)}.$
If the previous condition $\lambda(1)=\Gamma_{s}^{(h)}<\Gamma_{j}^{(i)}$ does not hold, we add the transition
(9) $(\Gamma_{j}^{(i)},a)\smash{\stackrel{{\scriptstyle\text{\tiny{$\lambda|\lambda^{\prime}\lambda$}\,}}}{{\relbar\joinrel\relbar\joinrel\longrightarrow}}}E,\lambda^{\prime}=(\Gamma_{j}^{(i)})\mbox{ whether }\lambda\mbox{ is final or not}.$
We will prove that each state $(\Gamma_{0},a)$ represents the action of the letter $a\in A$ of each vertex of $\Gamma$. We have the following lemma.
Lemma 8.3.
For any word $w=\lambda_{0}\ldots\lambda_{k}\in\Lambda^{*}$ formed by non-final letters, with $\lambda_{0}=(\Gamma_{0})$ and satisfying $\lambda_{0}<\ldots<\lambda_{k}$ we have:
$(\Gamma_{0},a)\ast w=\lambda_{0}\ldots\lambda_{k-1},\quad(\Gamma_{0},a)\cdot w=(\lambda_{k}(1),a)$
for any letter $a\in A$.
Proof.
It follows by induction on the length of $w$, where the base of induction is a consequence of (4) while the general case uses the transitions described in (5). ∎
Lemma 8.4.
For any infinite word $\eta\in\Lambda^{\omega}\setminus\left(\mathcal{F}(\Lambda)\Lambda^{\omega}\right)$ that is not well-formed, we have $(\Gamma_{0},a)\ast\eta=\eta$ for any letter $a\in A$.
Proof.
Let $w=\lambda_{0}\ldots\lambda_{k}$ be a (possibly empty) maximal prefix of $\eta$ formed by non-final letters, with $\lambda_{0}=(\Gamma_{0})$ and $\lambda_{0}<\ldots<\lambda_{k}$. If $w$ is empty, this means that $\lambda_{0}\neq(\Gamma_{0}),(\Gamma_{0},x_{0})$, hence by the transition described in (4) we deduce $(\Gamma_{0},a)\ast\eta=\eta$. Otherwise, $w$ is non-empty, and if $k$ is finite, then $\lambda_{k}\nless\lambda_{k+1}$. By Lemma 8.3 we have $(\Gamma_{0},a)\ast w=\lambda_{0}\ldots\lambda_{k-1}$ and $(\Gamma_{0},a)\cdot w=(\lambda_{k}(1),a)$. Hence, using the condition $\lambda_{k}\nless\lambda_{k+1}$ and by transition (9) we conclude that $(\Gamma_{0},a)\ast w\lambda_{k+1}=w\lambda_{k+1}$ and $(\Gamma_{0},a)\cdot w\lambda_{k+1}=E$, i.e. $(\Gamma_{0},a)\ast\eta=\eta$. If $k$ is infinite, meaning $w=\eta$, by Lemma 8.3 we may conclude our claim $(\Gamma_{0},a)\ast\eta=\eta$. ∎
We recall that for a vertex $x\in\Gamma$ and a letter $a\in A$ we denote by $x\cdot a$ the vertex such that $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}(x\cdot a)$ is an edge in $\Gamma$. We have the following lemma.
Lemma 8.5.
Let $a\in A$ and $x\in V(\Gamma)$. Then, $(\Gamma_{0},a)\ast w[x]\eta=w[x\cdot a]\eta$ for all $\eta\in\Lambda^{\omega}$.
Proof.
Let
$w[x]=\lambda_{0}\ldots\lambda_{\ell}=(\Gamma_{0})(\Gamma_{j_{1}}^{(i_{1})})\ldots(\Gamma_{j_{s}}^{(i_{s})})\ldots(\Gamma_{j_{\ell-1}}^{(i_{\ell-1})})(\Gamma_{j_{\ell}}^{(i_{\ell})},v_{\ell})$
be the well-formed word associated to the (unique) rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})$ with $x_{\ell}=x$, for which there are end-isomorphisms $\varphi_{s}:\Gamma(x_{s},x_{0})\to\Gamma_{j_{s}}^{(i_{s})}$ with $\varphi_{\ell}(x_{\ell})=v_{\ell}$ and satisfying $(\Psi(i_{s-1},j_{s-1})\ast\left.\varphi_{s-1})\right|_{\Gamma(x_{s},x_{0})}=\varphi_{s}$ for $s\in[1,\ell]$. By Lemma 8.3 we have
(10) $(\Gamma_{0},a)\ast\lambda_{0}\ldots\lambda_{\ell-1}=\lambda_{0}\ldots\lambda_{\ell-2},\quad(\Gamma_{0},a)\cdot w=(\lambda_{\ell-1}(1),a)$
Now, since $\lambda_{\ell}$ is final and $\lambda_{\ell-1}<\lambda_{\ell}$ we are in one of the three possibile transitions (6), (7), (8). Suppose that $\lambda_{\ell}=\Gamma_{s}^{(h)}$, and $\lambda_{\ell-1}=\Gamma_{j}^{(i)}$, where $\Gamma_{s}^{(h)}$ is a second-level end-cone of $\Psi(i,j)(\Gamma_{i}^{(j)})=\Gamma_{h}$.
  • Suppose that condition (6) holds. In this case $v_{\ell}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\Psi(i,j)(z)$ is an edge of $\Gamma_{h}$ with $z\in\Delta_{j}^{(i)}$, and so by (10) and by the definition we get:
    $(\Gamma_{0},a)\ast w\eta=\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)\eta,\quad\forall\eta\in\Lambda^{\omega}$
    We claim that $\Xi(x\cdot a)=\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)\Lambda^{\omega}$. Now, since $\Psi(i,j)^{-1}(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j}^{(i)}$ and $\varphi_{\ell}(x)=\varphi_{\ell-1}(x)$, we deduce that also $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell-1}(z)$ is an edge in $\Gamma(x_{\ell-1},x_{0})$ with $x\cdot a=\varphi^{-1}_{\ell-1}(z)\in\Delta(x_{\ell-1},x_{0})$. Therefore, by Lemma 8.1, the well-formed word $\lambda_{0}\ldots\lambda_{\ell-2}(\Gamma_{j}^{(i)},z)$ is associated the rooted end-geodesic:
    $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell-2},x_{0})\supset\Gamma(x\cdot a,x_{0})$
    and this concludes the proof of our claim.
  • Suppose that condition (7) holds. In this case we have
    $(\Gamma_{0},a)\ast w\eta=\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})},z)\eta,\quad\forall\eta\in\Lambda^{\omega}$
    where $v_{\ell}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge of $\Gamma_{j_{\ell}}^{(i_{\ell})}$ with $z\in\Delta_{j_{\ell}}^{(i_{\ell})}$. Now, since $\varphi_{\ell}(v_{\ell})=x$ and $\varphi^{-1}_{\ell}(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\varphi^{-1}_{\ell}(z)$ is an edge in $\Gamma(x_{\ell},x_{0})$, we conclude that $x\cdot a=\varphi^{-1}_{\ell}(z)$. Hence, $\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})},z)$ is the well-formed word associated to the rooted end-geodesic $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell-1},x_{0})\supset\Gamma(x\cdot a,x_{0})$, i.e., $(\Gamma_{0},a)\ast w[x]\eta=w[x\cdot a]\eta$ for all $\eta\in\Lambda^{\omega}$.
  • Suppose that condition (8) holds. Using the definition, we have
    $(\Gamma_{0},a)\ast w\eta=\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})})(\Gamma_{t}^{(m)},z)\eta,\quad\forall\eta\in\Lambda^{\omega}$
    where $\Gamma_{t}^{(m)}<\Gamma_{s}^{(h)}=\Gamma_{j_{\ell}}^{(i_{\ell})}$ and $\Psi(s,h)(v_{\ell})\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z$ is an edge in $\Gamma_{m}$ with $z\in\Delta_{t}^{(m)}$. Since $\Gamma_{t}^{(m)}$ is a second-level end-cone of $\Gamma_{m}$ and $\phi=\Psi(s,h)\circ\varphi_{\ell}:\Gamma(x,x_{0})\to\Gamma_{m}$ is an end-isomorphism, we deduce that $\phi^{-1}(\Gamma_{t}^{(m)})$ is a second level end-cone of $\Gamma(x,x_{0})$ containing $\phi^{-1}(z)$, and since $x\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\phi^{-1}(z)$ is an edge in $\Gamma(x,x_{0})$ we have $\phi^{-1}(\Gamma_{t}^{(m)})=\Gamma(x\cdot a,x_{0})$. Take $\varphi_{\ell+1}=\left.\phi\right|_{\Gamma(x\cdot a,x_{0})}$. We clearly have $(\left.\Psi(s,h)\circ\varphi_{\ell})\right|_{\Gamma(x\cdot a,x_{0})}=\varphi_{\ell+1}$. Thus, by Lemma 8.1 we have that the rooted end-geodesic
    $\Gamma(x_{0},x_{0})\supset\Gamma(x_{1},x_{0})\ldots\supset\Gamma(x_{\ell},x_{0})\supset\Gamma(x\cdot a,x_{0})$
    is associated to the well-formed word $\lambda_{0}\ldots\lambda_{\ell-1}(\Gamma_{j_{\ell}}^{(i_{\ell})})(\Gamma_{t}^{(m)},z)$ and this is our claim $(\Gamma_{0},a)\ast w[x]\eta=w[x\cdot a]\eta$ for all $\eta\in\Lambda^{\omega}$.
Theorem 8.6.
The transition group $\mathcal{G}(\Gamma^{(1)},\ldots,\Gamma^{(k)})$ of a collection of context-free complete graphs $\Gamma^{(1)}\ldots,\Gamma^{(k)}$ is rational.
Proof.
Since rationality is closed by taking direct products and pass to subgroups, by Lemma 4.4 we reduce to the case $\mathcal{G}(\Gamma)$.
As a shorthand notation let us put $\gamma_{a}=(\Gamma_{0},a)$, $a\in A$. By Lemma 8.5 and the fact that $\Gamma$ is an inverse graph, we have $(\gamma_{a^{-1}}\gamma_{a})\ast w[x]\eta=w[x]\eta$ and $(\gamma_{a}\gamma_{a^{-1}})\ast w[x]\eta=w[x]\eta$ for all $\eta\in\Lambda^{\omega}$. Moreover, by Lemma 8.4 for all $\xi\in\Lambda^{\omega}\setminus\mathcal{F}(\Lambda)$ we have $(\gamma_{a^{-1}}\gamma_{a})\ast\xi=\xi$ and $(\gamma_{a}\gamma_{a^{-1}})\ast\xi=\xi$. Therefore, we may conclude that each state $\gamma_{a}$ of $\mathcal{T}_{\Gamma}$ gives rise to a rational bijection $\gamma_{a}\ast:\Lambda^{\omega}\to\Lambda^{\omega}$ with inverse $\gamma_{a^{-1}}\ast$. Thus, by [3, Proposition 1.3], each state $\gamma_{a}$ is a rational homeomorphism of $\Lambda^{\omega}$. Let us prove that the subgroup $H(\Gamma)=\langle\gamma_{a}:a\in A\rangle$ of the rational group $\mathcal{R}_{\Lambda}$ of all the rational homeomorphisms of $\Lambda^{\omega}$ is isomorphic to the transtion group $\mathcal{G}(\Gamma)$. Consider a word $u=a_{1}\ldots a_{p}\in A^{*}$. By Lemma 8.5, and using a simple induction on the length $p$ of a word $u$ we have that the following equality
$(\gamma_{a_{k}}\ldots\gamma_{a_{1}})\ast w[x]\eta=w[x\cdot u]\eta,\,\forall\eta\in\Lambda^{\omega}$
holds for any vertex $x\in\Gamma$. It is clear that if $\gamma_{a_{k}}\ldots\gamma_{a_{1}}$ represents the identity in $H(\Gamma)$, then
$w[x]\eta=(\gamma_{a_{k}}\ldots\gamma_{a_{1}})\ast w[x]\eta=w[x\cdot u]\eta.$
Now, since the well-formed words form a prefix code, then for an infinite word $\eta$ there is at most one (maximal) prefix that is well-formed, by the previous equality we conclude that $w[x]=w[x\cdot u]$ for all $x\in V(\Gamma)$, i.e., $u\in\mathcal{L}(\Gamma)$ represents the identity in $\mathcal{G}(\Gamma)$. Therefore, the isomorphism $\phi:\{\gamma_{a}:a\in A\}^{*}\to A$ defined by $\phi(\gamma_{a})=a$ extends to the epimomorphism $\varphi:H(\Gamma)\to\mathcal{G}(\Gamma)$ naturally. To prove that $\varphi$ is an isomorphism suppose that $\varphi(\gamma_{a_{k}}\ldots\gamma_{a_{1}})=u$ represents the identity in $\mathcal{G}(\Gamma)$, then $x\cdot u=x$ for all $x\in\Gamma$, that is equivalent to $w[x]=w[x\cdot u]$ for all $x\in\Gamma$ which in turn is equivalent to $(\gamma_{a_{k}}\ldots\gamma_{a_{1}})\ast w[x]\eta=w[x]\eta$ for all $\eta\in\Lambda^{\omega}$, $x\in V(\Gamma)$. Now, if $\xi$ is not well-formed, by Lemma 8.4 we have $(\gamma_{a_{k}}\ldots\gamma_{a_{1}})\ast\xi=\xi$. Therefore, we have shown that $(\gamma_{a_{k}}\ldots\gamma_{a_{1}})\ast\xi=\xi$ for all $\xi\in\Lambda^{\omega}$, i.e., $\gamma_{a_{k}}\ldots\gamma_{a_{1}}$ represents the identity in $H(\Gamma)$. This concludes the proof that $H(\Gamma)\simeq\mathcal{G}(\Gamma)$, i.e., $\mathcal{G}(\Gamma)$ embeds into the rational group $\mathcal{R}_{\Lambda}$. ∎
Note that, despite we already have a tree structure associated to the graph, namely the tree of end-cones, the transducer acts on $\Lambda^{\omega}$ which contains a copy of such tree as a subtree.
Remark 8.7.
The transducer that represents any element $g$ in a CF-TR group is bi-lipschitz. It suffices to notice that any element in $\mathcal{F}(\Lambda)$ is shortened or lengthened by at most $|g|$ letters.
9. Free product of context-free inverse graphs
The aim of this section is to introduce the operation of free product of graphs and proving that the property of being context-free is preserved under this operation. This is useful in constructing examples of context-free graphs from simpler ones.
Figure 2. A depiction of the free product of two rooted graphs $(\Gamma,\gamma_{0}),(\Lambda,\lambda_{0})$ with gluing maps $\psi^{1}(v)=\lambda_{0}$ for all $v\in\Gamma\setminus\{\gamma_{0}\}$ and $\psi^{2}(v)=\gamma_{0}$ for all $v\in\Lambda\setminus\{\lambda_{0}\}$.
Definition 9.1 (Free product of graphs).
Let $\Theta_{1}=\{(\Gamma^{(i)},\gamma^{i}_{0}),i=1,\ldots,k\}$ be a collection of $A_{1}$-digraphs, and similarly let $\Theta_{2}=\{(\Lambda^{(j)},\lambda^{(j)}_{0}),j=1,\ldots,m\}$ be a collection of $A_{2}$-digraphs, with $A_{1}\cap A_{2}=\emptyset$. For each $i\in[1,k]$, $j\in[1,m]$ we consider the functions $\psi^{1}_{i}:V(\Gamma^{(i)})\setminus\{\gamma_{0}^{i}\}\to\{\lambda^{1}_{0},\ldots,\lambda^{m}_{0}\}$ and $\psi^{2}_{j}:V(\Lambda^{(j)})\setminus\{\lambda_{0}^{j}\}\to\{\gamma^{1}_{0},\ldots,\gamma^{k}_{0}\}$; the gluing map is the collection $\Psi$ of all such functions. We first consider the infinite disjoint union of copies of the graphs $\Theta_{1},\Theta_{2}$ where the set of vertices are words alternating between vertices of the graphs $\Theta_{1}$ and $\Theta_{2}$. More precisely, we consider the $A_{1}\cup A_{2}$-digraph $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ whose set of vertices $\mathcal{V}$ is formed by alternating sequences of vertices of $\Theta_{1},\Theta_{2}$ which are either of the form $\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{\ell}},\mbox{ or }\lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{\ell}}$, with $\gamma^{i_{n}}_{s_{n}}\in V(\Gamma^{(i_{n})}),\lambda^{j_{n}}_{t_{n}}\in V(\Lambda^{(j_{n})})$, $n\in[1,\ell]$ and $\ell\in\mathbb{N}$. The edges are of the form
$\begin{cases}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{2}}_{s_{2}}\lambda^{j_{2}}_{t_{2}}\ldots\gamma^{i_{\ell}}_{s_{\ell}}\lambda^{j_{\ell}}_{t_{h}},a\in A_{2},\mbox{ whenever }\lambda^{j_{\ell}}_{t_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{\ell}}_{t_{h}}\in E(\Lambda^{(j_{\ell})})\\ \lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda^{j_{1}}_{t_{1}}\gamma^{i_{1}}_{s_{1}}\lambda^{j_{2}}_{t_{2}}\gamma^{i_{2}}_{s_{2}}\ldots\lambda^{j_{\ell}}_{t_{\ell}}\gamma^{i_{\ell}}_{s_{h}},a\in A_{1},\mbox{ whenever }\gamma^{j_{\ell}}_{s_{\ell}}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma^{j_{\ell}}_{s_{h}}\in E(\Gamma^{(j_{\ell})})\end{cases}$
Roughly speaking we are taking infinitely many copies of the graphs $\Gamma^{(i)},\Lambda^{(j)}$ where we simply rename each vertex by adding a prefix that is a word that is alternating in the sense explained above. Now, the free product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is the $A_{1}\cup A_{2}$-digraph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the transitive closure of the following relation:
$\gamma_{0}^{1}v^{*}=\lambda_{0}^{1}v^{*},\;v^{*}\gamma_{i}^{j}u^{*}=v^{*}\psi_{j}^{1}(\gamma_{i}^{j})u^{*},\;v^{*}\lambda_{i}^{j}u^{*}=v^{*}\psi_{j}^{2}(\lambda_{i}^{j})u^{*}$
where $u^{*},v^{*}$ are generic alternating words on $\Theta_{1},\Theta_{2}$ possibly empty. Henceforth, we will consider the rooted graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ where $\gamma_{0}^{1}=\lambda_{0}^{1}$.
Roughly speaking the rooted graph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is obtained from the graphs $\Gamma^{(1)}\sqcup\Lambda^{(1)}$ by first identifying $\gamma_{0}^{1}=\lambda_{0}^{1}$, and then by iteratively gluing at a vertex with representative $v^{*}\gamma^{j}_{i}$, with $i\neq 0$ (not a root), the graph $(\Lambda^{(s)},\lambda^{s}_{0})$ with $\lambda^{s}_{0}=\psi^{1}_{j}(\gamma^{j}_{i})$, so that the two vertices $v^{*}\gamma_{i}^{j},v^{*}\lambda_{0}^{s}$ identify. In case the vertex ends with a vertex from the $\Lambda$ graphs such as $v^{*}\lambda^{j}_{i}$, then we glue the graph $(\Gamma^{(s)},\gamma_{0}^{s})$ by making the identification $\psi^{2}_{j}(\lambda^{j}_{i})=\gamma_{0}^{s}$. Let us call a vertex $v^{*}\gamma^{j}_{i}$, where the last vertex belongs to some graph of the collection $\Theta_{1}$, a vertex colored by $1$, while if it ends by a vertex from the collection $\Theta_{2}$ we say that it is colored by $2$. To have an idea of what kind of digraph we may generate see Fig. 2 in which the free product of two rooted graphs $(\Gamma,\gamma_{0}),(\Lambda,\lambda_{0})$ is depicted, where the gluing maps are $\psi^{1}(v)=\lambda_{0}$ for all $v\in\Gamma\setminus\{\gamma_{0}\}$ and $\psi^{2}(v)=\gamma_{0}$ for all $v\in\Lambda\setminus\{\lambda_{0}\}$. We will see later that if $\Gamma,\Lambda$ are two Cayley graphs of two groups $G_{1},G_{2}$, then $(\Gamma,\gamma_{0})\star_{\Psi}(\Lambda,\lambda_{0})$ is the Cayley graph of the free product $G_{1}\star G_{2}$, hence the name. The following proposition is an easy consequence of the definition and the condition $A_{1}\cap A_{2}=\emptyset$.
Proposition 9.2.
If $\Theta_{1}$ ($\Theta_{2}$) is a family of deterministic (complete) $A_{1}$-digraphs ($A_{2}$-digraphs), then $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a deterministic (complete) $A_{1}\cup A_{2}$-digraph. In particular, if $\Theta_{1}$ ($\Theta_{2}$) is a family of inverse $A_{1}$-graphs ($A_{2}$-graphs), then the rooted graphs $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$ is an inverse $A_{1}\cup A_{2}$-graph.
We now make some geometric considerations on this construction. For simplicity of notation let $\alpha=\gamma_{0}^{1}=\lambda_{0}^{1}$ and put $\Omega=\Theta_{1}\star_{\Psi}\Theta_{2}$. Now, consider a geodesic $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}$ connecting $\alpha$ to some vertex $v^{*}\gamma_{i}^{j}$ of $(\Omega,\alpha)$, colored by $1$ with $i\neq 0$. Since $i\neq 0$, $w$ ends with a word colored by $1$, thus it may be uniquely factor as an alternating product $w=z_{1}x_{1}z_{2}x_{2}\ldots z_{k}x_{k}$ with $z_{t}\in A_{2}$ and $x_{t}\in A_{1}$. Let $\Lambda^{(s)}$ be the digraph of $\Theta_{2}$ with $\lambda_{0}^{s}=\psi^{1}_{j}(\gamma_{i}^{j})$ so that in $\Omega$, the vertex $v^{*}\gamma_{i}^{j}$ is identified with $v^{*}\lambda_{0}^{s}$. Note that we are gluing at $v^{*}\gamma_{i}^{j}$ the rooted $A_{1}\cup A_{2}$-digraph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ obtained from $\Lambda^{(s)}$ in which we iteratively glue all the digraphs from the collection $\Theta_{1},\Theta_{2}$ according to the gluing maps $\Psi$. More formally, this subgraph is isomorphic to the $A_{1}\cup A_{2}$-digraph $\Lambda^{(s)}(\uparrow_{\Psi})$ which is the connected component containing $\lambda_{0}^{s}$ of the digraph obtained from $(\Theta_{1}\sqcup\Theta_{2})^{\oplus}$ by identifying the vertices of $\mathcal{V}$ via the equivalence relation generated by the relation
$v^{*}\gamma_{i}^{j}u^{*}=v^{*}\psi_{j}^{1}(\gamma_{i}^{j})u^{*},\;v^{*}\lambda_{i}^{j}u^{*}=v^{*}\psi_{j}^{2}(\lambda_{i}^{j})u^{*}$
Note that $\Lambda^{(s)}(\uparrow_{\Psi})\simeq\Omega^{+}(v^{*}\gamma_{i}^{j})$ via the map on the set of vertices defined by $\phi(\lambda_{0}^{s})=v^{*}\gamma_{i}^{j}=v^{*}\lambda_{0}^{s}$ and $\phi(\lambda_{0}^{s}u^{*})=v^{*}\lambda_{0}^{s}u^{*}$. The digraph $\Omega^{+}(v^{*}\gamma_{i}^{j})$ is called the out-cone of $\Omega$ at $v^{*}\gamma_{i}^{j}$. Moreover, since we are gluing a copy of $\Lambda^{(s)}(\uparrow_{\Psi})$ by identifying the vertex $\lambda_{0}^{s}$ with $v^{*}\gamma_{i}^{j}$, this means that all the geodesics $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$ of $\Omega$ passing through $v^{*}\gamma_{i}^{j}$ with $w^{\prime}=u_{1}u_{2}$, for some $u_{1}\in A_{2}^{+}$, have the property that $v^{*}\gamma_{i}^{j}u^{*}$ belongs to the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, see Fig. 3 for a graphical representation. We record this remark in the following lemma.
Lemma 9.3.
Put $(\Omega,\alpha)=(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma_{0}^{1})$, then for any $\gamma_{i}^{j},\lambda_{i}^{j}$ with $i\neq 0$ we have:
$V(\Omega^{+}(v^{*}\gamma_{i}^{j}))=\{v^{*}\gamma_{i}^{j}\}\cup\{v^{*}\gamma_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{2}^{+}(A_{1}\cup A_{2})^{*}\}$
$V(\Omega^{+}(v^{*}\lambda_{i}^{j}))=\{v^{*}\lambda_{i}^{j}\}\cup\{v^{*}\lambda_{i}^{j}u^{*}:\mbox{ there is a geodesic }\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\lambda_{i}^{j}u^{*}\mbox{ with }w^{\prime}\in A_{1}^{+}(A_{1}\cup A_{2})^{*}\}$
Moreover, for any vertex $v^{*}\gamma_{i}^{j}=v^{*}\lambda_{0}^{s}$ colored by $1$, or $v^{*}\lambda_{i}^{j}=v^{*}\gamma_{0}^{s}$ colored by $2$, we have:
$\Omega^{+}(v^{*}\gamma_{i}^{j})\simeq\Lambda^{(s)}(\uparrow_{\Psi}),\quad\Omega^{+}(v^{*}\lambda_{i}^{j})\simeq\Gamma^{(h)}(\uparrow_{\Psi}).$
Figure 3. In the dashed region the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})$, in red the geodesic $\alpha\smash{\stackrel{{\scriptstyle\text{\tiny{$w$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$w^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}v^{*}\gamma_{i}^{j}u^{*}$.
We now describe how to obtain the end-cones of $\Omega$ from the end-cones of the graphs in $\Theta_{1}\cup\Theta_{2}$. We have the following lemma which, for simplicity, we state for a generic vertex colored by $2$ but it clearly holds for any vertex of $\Omega$.
Lemma 9.4.
With the above notation, for a given vertex $v^{*}\gamma_{i}^{j}\lambda^{s}_{t}$ of $\Omega$ colored by $2$, the end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ is isomorphic to the $A_{1}\cup A_{2}$-digraph obtained by taking the disjoint union
$\Lambda^{(s)}(\lambda^{s}_{t},\lambda^{s}_{0})\cup\biguplus_{\begin{subarray}{c}z\in V(\Lambda^{(s)}(\lambda^{s}_{t},\lambda^{s}_{0}))\\ \gamma^{h}_{0}=\psi^{2}_{s}(z)\end{subarray}}\Gamma^{(h)}(\uparrow_{\Psi})$
and then by identifying each vertex $z\in V(\Lambda^{(s)}(\lambda^{s}_{t},\lambda^{s}_{0}))$ with $\gamma^{h}_{0}=\psi^{2}_{s}(z)$.
Proof.
Set $n=\|v^{*}\gamma_{i}^{j}\lambda^{s}_{t}\|_{\alpha}$. Then, by Lemma 9.3 we may conclude that the end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ is contained in the out-cone $\Omega^{+}(v^{*}\gamma_{i}^{j})\simeq\Lambda^{(s)}(\uparrow_{\Psi})$. Moreover, the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ are in one to one correspondence with the frontier vertices of $\Lambda^{(s)}(\lambda_{t}^{s},\lambda_{0}^{s})$. Therefore, $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ may be obtained from $\Lambda^{(s)}(\lambda_{t}^{s},\lambda_{0}^{s})$ by gluing at each vertex $v$ of $\Lambda^{(s)}(\lambda_{t}^{s},\lambda_{0}^{s})$ the corresponding out-cone $\Gamma^{(h)}(\uparrow_{\Psi})$ by identifying $z$ with $\gamma_{0}^{h}=\psi^{2}_{s}(z)$. See Figure 4 for a graphical representation on how to obtain $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ from $\Lambda^{(s)}(\lambda_{t}^{s},\lambda_{0}^{s})$. ∎
Figure 4. A depiction of the end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda^{s}_{t},\alpha)$ which is obtained by gluing at each vertex $z$ of the end-cone $\Lambda^{(s)}(\lambda^{s}_{t},\lambda^{s}_{0})$ of $\Lambda^{(s)}$ (in dashed red lines) the out-cone $\Gamma^{(h)}(\uparrow_{\Psi})$ according to the value of gluing map $\gamma_{0}^{h}=\psi_{s}^{2}(z)$.
9.1. Context-free graphs and free product
In this section we prove that the free product of directed graphs preserves the property of being context-free. For this purpose, we need to slightly adapt the notion of isomorphism and end-cone of an $A_{1}$-digraph to include the notion of vertex coloring. A vertex-coloring of a graph $\Gamma$ is just a map $\chi:V(\Gamma)\to C$ onto some set of colors $C$. An isomorphism $\psi:\Gamma^{(1)}\to\Gamma^{(2)}$ between two $A$-digraphs with color maps $\chi_{1}:V(\Gamma)\to C$, $\chi_{2}:V(\Gamma^{(2)})\to C$ is an isomorphism of $A$-digraphs that preserves the colorings: $\chi_{1}=\chi_{2}\circ\psi$. For technical reasons that will be clear later, we call a coloring of a rooted $A$-digraph $(\Gamma,\gamma_{0})$ any map $\chi:V(\Gamma)\setminus\{\gamma_{0}\}\to C$, and we use the notation $(\Gamma,\gamma_{0};\chi)$ to stress the coloring function. For such colored rooted $A$-digraph $(\Gamma,\gamma_{0};\chi)$, we may regard an end-cone $\Gamma(v,\gamma_{0})$ as a colored $A$-digraph with coloring map $\chi\mid_{V(\Gamma(v,\gamma_{0}))}$ given by the restriction of $\chi$ to the set of vertices of the end-cone $\Gamma(v,\gamma_{0})$. We therefore say that the colored $A$-digraph $(\Gamma,\gamma_{0};\chi)$ is context-free if there are finitely many end-cones up to end-isomorphism (in the aforementioned sense). We stick to the notation introduced in Definition 9.1, and let us suppose that $\Theta_{1}$ ($\Theta_{2}$) is a collection of $A_{1}$-digraphs ($A_{2}$-digraphs) with $A_{1}\cap A_{2}=\emptyset$ and with gluing maps $\Psi$. We say that $\Gamma^{(j)}\in\Theta_{1}$ ($\Lambda^{(j)}\in\Theta_{2}$) is context-free if the colored rooted $A_{1}$-digraph $(\Gamma^{(j)},\gamma_{0}^{j},\psi^{1}_{j})$ ($A_{2}$-digraph $(\Lambda^{(j)},\lambda_{0}^{j},\psi^{2}_{j})$) is context-free. If all the colored $A_{i}$-digraphs of $\Theta_{1}\cup\Theta_{2}$ are context-free, we say that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble. We have the following theorem.
Theorem 9.5.
For a context-free ensemble $(\Theta_{1},\Theta_{2},\Psi)$, the free-product $\Theta_{1}\star_{\Psi}\Theta_{2}$ is a context-free $A_{1}\cup A_{2}$-digraph.
Proof.
By Lemma 9.4, using the same notation, we have that an end-cone $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ is obtained from the disjoint union
$\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0})\cup\biguplus_{\begin{subarray}{c}z\in V(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))\\ \gamma^{h}_{0}=\psi^{2}_{s}(z)\end{subarray}}\Gamma^{(h)}(\uparrow_{\Psi})$
an then by identifying $z$ with $\psi^{2}_{s}(z)$. We prove that we have finitely many end-cones up to end-isomorphism (in the usual sense). Now, let $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$ be another end-cone for which there is an end-isomorphism $\varphi:\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0})\to\Lambda^{s}(\lambda^{s}_{\ell},\lambda^{s}_{0})$ of end-cones of the same $A_{2}$-digraph $\Lambda^{(s)}$ that preserves the coloring. Since the coloring map $\psi^{2}_{s}$ is commuting with $\varphi$, for any $z\in V(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$ we have that $\psi^{2}_{s}(z)=\psi^{2}_{s}(\varphi(z))=\gamma_{0}^{h}$, so both $z$ and $\varphi(z)$ have the same color, hence in both the vertex $z$ and $\varphi(z)$ we are gluing the same digraph $\Gamma^{(h)}(\uparrow_{\Psi})$. Therefore, it is straightforward to show that $\varphi$ may be extended
$\widehat{\varphi}:\left(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0})\cup\biguplus_{\begin{subarray}{c}z\in V(\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))\\ \gamma^{h}_{0}=\psi^{2}_{s}(z)\end{subarray}}\Gamma^{(h)}(\uparrow_{\Psi})\right)\to\left(\Lambda^{s}(\lambda^{s}_{\ell},\lambda^{s}_{0})\cup\biguplus_{\begin{subarray}{c}z\in V(\Lambda^{s}(\lambda^{s}_{\ell},\lambda^{s}_{0}))\\ \gamma^{h}_{0}=\psi^{2}_{s}(z)\end{subarray}}\Gamma^{(h)}(\uparrow_{\Psi})\right)$
by locally extending $\varphi$ from each disjoint copy $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $z$, to the same digraph $\Gamma^{(h)}(\uparrow_{\Psi})$ associated to $\varphi(z)$. Now, since we are performing the identification $z=\psi^{2}_{s}(z)$ in the domain of $\widehat{\varphi}$ and this is the same as identifying $\varphi(z)$ with $\psi^{2}_{s}(\varphi(z))$, we have that $\widehat{\varphi}$ induces an isomorphism $\phi:\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)\to\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Note that the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ are (up to renaming) the same as $\Lambda^{s}(\lambda^{s}_{t},\lambda^{s}_{0}))$, and similarly for $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$. Hence, $\phi$ is mapping the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{t}^{s},\alpha)$ into the frontier vertices of $\Omega(v^{*}\gamma_{i}^{j}\lambda_{\ell}^{s},\alpha)$, hence they are end-isomorphic. We have just shown that the end-isomorphism classes of the end-cones of $\Omega$ of vertices colored by $2$ depends on the end-isomorphism (preserving the coloring) classes of the digraphs $\{\Lambda^{(1)},\ldots,\Lambda^{(m)}\}$ of $\Theta_{2}$, which are finitely many. Similarly, one may show the same for the end-cones of the vertices colored by $1$, and this concludes the proof of the theorem. ∎
Question 9.6.
It would be interesting to see what is the corresponding construction from the machine point of view. More precisely, if one considers the pushdown automata $M_{i}$ that accepts the language of the words labelling a circuit at each root of each context-free graph of $\Theta_{1}\cup\Theta_{2}$, what is the construction, involving each $M_{i}$, that generates the pushdown-automaton accepting the language of circuit of the digraph $\Theta_{1}\star_{\Psi}\Theta_{2}$?
Figure 5. On the upper part the digraphs $\Theta_{1},\Theta_{2}$, in the lower picture the antenna digraph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
10. Some examples using the free product of inverse graphs
In this section we use the free product of context-free graphs to construct some examples of interesting CF-TR groups. In the first example, the antenna digraph, we show that in general transition groups of context-free graphs are not residually-finite. In the second one, the comb digraph, we prove that the associated transition group is a co-context-free group that is not poly-context-free, showing that these classes of groups are different. The third example addresses the question about the possibility of having arbitrarily large finite subgroups, we answer positively to this by providing a infinite torsion free subgroup.
Definition 10.1 (The Antenna digraph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1},b,b^{-1}\}$. Consider the rooted inverse $A_{1}$-digraph $(\Gamma,\gamma)$ which is the line
$\gamma_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$c$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i+1},\;\gamma_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$c^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\gamma_{i},\quad\forall i\in\mathbb{Z},\mbox{ where the root }\gamma=\gamma_{0}$
and the two rooted inverse $A_{2}$-digraphs $(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})$ which are lines with loops defined by
$\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\;\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\,\lambda_{a,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i+1},\;\lambda_{a,i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{a,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{a,0}=\lambda_{a}$
$\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\;\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\,\lambda_{b,i}\smash{\stackrel{{\scriptstyle\text{\tiny{$b$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i+1},\;\lambda_{b,i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$b^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}\lambda_{b,i},\quad\forall i\in\mathbb{Z},\,\,\lambda_{b,0}=\lambda_{b}$
see the upper part of Fig. 5. The gluing map $\Psi$ is given by the functions defined as follows:
$\psi^{1}:V(\Gamma)\setminus\{\gamma\}\to\{\lambda_{a},\lambda_{b}\},\,\,\psi^{1}(\gamma_{i})=\begin{cases}\lambda_{b}&\mbox{ if }i\leq 0\\ \lambda_{a}&\mbox{ otherwise}\end{cases}$
and
$\psi^{2}_{a}:V(\Lambda_{a})\setminus\{\lambda_{a}\}\to\{\gamma\},\,\,\psi^{2}_{b}:V(\Lambda_{b})\setminus\{\lambda_{b}\}\to\{\gamma\}$
that are constant maps. By putting $\Theta_{1}=\{(\Gamma,\gamma)\}$ and $\Theta_{2}=\{(\Lambda_{a},\lambda_{a}),(\Lambda_{b},\lambda_{b})\}$, the rooted inverse $A_{1}\cup A_{2}$-digraph $(\mathfrak{A},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$ is called the antenna digraph, see Fig. 5. Moreover, it is easy to see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free-ensemble, thus by Theorem 9.5 $(\mathfrak{A},\gamma)$ is a context-free inverse graph. Note that with the convention of the free product that we have given, the root $\gamma$ of $\Theta_{1}\star_{\Psi}\Theta_{2}$ is identified with $\lambda_{b}$.
By the tree structure, the set of vertices can be identified using words in $A_{1}\cup A_{2}$ by associating to a vertex the unique geodesic connecting $\gamma$ to such vertex. It is easy to see that the action of the alphabet $A_{1}\cup A_{2}$ on the set of vertices is defined as follows. Suppose we are in the vertex $v=wx^{s}$, $s\neq 0$, in case $x\in A_{1}\cup A_{2}$ we have
(11) $wc^{s}\cdot b^{t}=\begin{cases}wc^{s},&\mbox{ if }s>0\\ wc^{s}b^{t},&\mbox{ otherwise }\end{cases},\,\,wc^{s}\cdot a^{t}=\begin{cases}wc^{s}a^{t},&\mbox{ if }s\geq 0\\ wc^{s},&\mbox{ otherwise }\end{cases}\,\,$
$wa^{s}\cdot b^{t}=wa^{s},\,\,wb^{s}\cdot a^{t}=wb^{s},\,\,wa^{s}\cdot c^{t}=wa^{s}c^{t},\,\,wb^{s}\cdot c^{t}=wb^{s}c^{t},\,\,\gamma\cdot a^{s}=\gamma,\gamma\cdot b^{s}=\gamma b^{s}$
We now prove that $\mathcal{G}(\mathfrak{A})$ is not residually finite, but first we need the following lemma.
Lemma 10.2.
For each $k,s\in\mathbb{Z}$, $t,p\geq 0$ we have that $\left[c^{-t}a^{s}c^{t},c^{p}b^{k}c^{-p}\right]=\mathds{1}_{\mathcal{G}(\mathfrak{A})}.$
Proof.
Note that the statement is equivalent to $g=\left[a^{s},c^{\alpha}b^{k}c^{-\alpha}\right]=a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha}=\mathds{1}_{\mathcal{G}(\mathfrak{A})}$ for every $\alpha\geq 0$ and $k,s\neq 0$. We may also assume that $\alpha>0$ since the relation $\left[a,b\right]=\mathds{1}_{\mathcal{G}(\mathfrak{A})}$ can be easily verified.
Note that $\gamma\cdot g=\gamma$. We show that for any vertex $v=wx^{m}$, $m\neq 0$ $x\in A_{1}\cup A_{2}$ we have $v\cdot g=v$ from which the statement follows. We have the following cases:
  • $m>0$, $x=c$, using (11) we have
    $wc^{m}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})$ $=wc^{m}a^{s}c^{\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wc^{m}a^{s}c^{\alpha}\cdot(c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})$
    $=wc^{m}\cdot(c^{\alpha}b^{-k}c^{-\alpha})=wc^{m},$
  • $m<0$, $x=c$, using (11) we have $wc^{m}\cdot g=wc^{{m+\alpha}}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})$. Now we have the following cases:
    • if $m+\alpha>0$ we have $wc^{m+\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wc^{m}\cdot(c^{\alpha}b^{-k}c^{-\alpha})=wc^{m}$.
    • if $m+\alpha<0$ we have $wc^{m+\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wc^{m+\alpha}b^{k}\cdot(c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wc^{m+\alpha}b^{k}c^{-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wc^{m+\alpha}b^{k}c^{-\alpha}\cdot(c^{\alpha}b^{-k}c^{-\alpha})=wc^{m}$.
    • if $m+\alpha=0$, by (11) we have that in the most general form, the vertex $v$ is either of the form $v=v^{\prime}c^{-\ell}b^{t}c^{-\alpha}$ or $v=v^{\prime}c^{\ell}a^{t}c^{-\alpha}$ for some $\ell>0$. Let us consider these two further cases:
        In the first case we have:
        $v^{\prime}c^{-\ell}b^{t}c^{-\alpha}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})$
        Now, if $t\neq-k$, then $v^{\prime}c^{-\ell}b^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t+k}c^{-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{t}c^{-\alpha}=v$. Otherwise, suppose that $t=-k$. In this case we have
        $v^{\prime}c^{-\ell}b^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
        $=v^{\prime}c^{-\ell}\cdot(b^{-k}c^{-\alpha})=v^{\prime}c^{-\ell}b^{-k}c^{-\alpha}=v.$
        Hence, $v\cdot g=v$.
        In the second case we the following computation:
        $v^{\prime}c^{\ell}a^{t}c^{-\alpha}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
        $=v^{\prime}c^{\ell}a^{t}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{\ell}a^{t}\cdot(c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
        $=v^{\prime}c^{\ell}a^{t}c^{-\alpha}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{\ell}a^{t}\cdot(b^{-k}c^{-\alpha})=v^{\prime}c^{\ell}a^{t}c^{-\alpha}=v$
  • if $x=b$ we have the following computation:
    $wb^{m}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wb^{m}c^{\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
    $=wb^{m}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wb^{m}c^{\alpha}\cdot(b^{-k}c^{-\alpha})=wb^{m}$
  • the last case is when $x=a$. In this case we have to subdivide the following two sub-cases:
    • if $m+s\neq 0$, for which we have:
      $wa^{m}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wa^{m+s}c^{\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
      $=wa^{m+s}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=wa^{m}\cdot(c^{\alpha}b^{-k}c^{-\alpha})=wa^{m}.$
[…]
      $v^{\prime}c^{t}a^{-s}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{t+\alpha}\cdot(b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=$
      $=v^{\prime}c^{t}\cdot(a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=v^{\prime}c^{t}a^{-s}$
      In case $t=0$ we necessarily have $v^{\prime}=\gamma$ and also in this case we have $\gamma a^{-s}\cdot(a^{s}c^{\alpha}b^{k}c^{-\alpha}a^{-s}c^{\alpha}b^{-k}c^{-\alpha})=\gamma a^{-s}$.
Theorem 10.3.
The group $G=\mathcal{G}(\mathfrak{A})$ defined by the antenna inverse graph $\mathfrak{A}$ is not residually finite.
Proof.
Let us assume that $G$ is residually finite. Consider the element
$h=\left[a,c^{-1}bc\right]=ac^{-1}bca^{-1}c^{-1}b^{-1}c$
Note that $h\neq\mathds{1}_{\mathcal{G}(\mathfrak{A})}$ since we have:
$\gamma\cdot ac^{-1}bca^{-1}c^{-1}b^{-1}c=\gamma\cdot(c^{-1}bca^{-1}c^{-1}b^{-1}c)=\gamma c^{-1}bca^{-1}c^{-1}b^{-1}c\neq\gamma$
Now, since $G$ is residually finite, there is a finite group $H$ and an homomorphism $\varphi:G\to H$ such that $\varphi(h)\neq\mathds{1}_{H}$. Since $H$ is finite there is some positive $\alpha>0$ such that $\varphi(c)^{\alpha}=\varphi(c)^{-1}$. By Lemma 10.2 we have $\left[a,c^{\alpha}bc^{-\alpha}\right]=\mathds{1}_{\mathcal{G}(\mathfrak{A})}$, hence:
$\mathds{1}_{H}$ $=\varphi\left(\left[a,c^{\alpha}bc^{-\alpha}\right]\right)=\left[\varphi(a),\varphi(c)^{\alpha}\varphi(b)\varphi(c)^{-\alpha}\right]=\left[\varphi(a),\varphi(c)^{-1}\varphi(b)\varphi(c)\right]=$
$=\varphi\left(\left[a,c^{-1}bc\right]\right)=\varphi(h)\neq\mathds{1}_{H}$
a contradiction, hence $G$ is not residually finite. ∎
In light of Theorem 6.2 and the preceding theorem, we observe that, if Brough Conjecture 1 regarding poly-context-free groups is true, then the class of all poly-context-free groups would be strictly contained within the class CF-TR. Although we are not able to show that the class of poly-context-free is included into our class, in the next example we show that these two classes are different. This is done by defining the following context-free digraph.
Figure 6. On the upper part the digraphs $\Theta_{1},\Theta_{2}$, in the lower picture the comb digraph $(\Theta_{1}\star_{\Psi}\Theta_{2},\gamma)$.
Definition 10.4 (The comb digraph).
Let $A_{1}=\{c,c^{-1}\}$, $A_{2}=\{a,a^{-1}\}$. Let us define the rooted inverse $A_{1}$-digraphs $(\Gamma,z_{0})$ whose underlying digraph is the line
$z_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$c$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i+1},\;z_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$c^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i},\quad\forall i\in\mathbb{Z},$
The other two rooted inverse $A_{2}$-digraphs $(\Lambda_{a},\lambda_{a}),(\Lambda,x_{0})$ are defined as follows. The inverse graph $(\Lambda_{a},\lambda_{a})$ consists of just one vertex $\lambda_{a}$ with loops labelled by $a,a^{-1}$. Instead, $(\Lambda_{a},x_{0})$ is the defined as the line on $A_{2}$:
$x_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i+1},\;x_{i+1}\smash{\stackrel{{\scriptstyle\text{\tiny{$a^{-1}$}}}}{{\relbar\joinrel\longrightarrow}}}x_{i},\quad\forall i\in\mathbb{Z},$
See the upper part of Fig. 6. The gluing map $\Psi$ is given by the following functions. The gluing maps are given by the following constant maps
$\psi_{1}:V(\Gamma)\setminus\{z_{0}\}\to\{\lambda_{a}\},\,\quad\psi_{1}(z)=\lambda_{a}$
$\psi_{2}:V(\Lambda)\setminus\{x_{0}\}\to\{z_{0}\},\,\quad\psi_{2}(x_{i})=z_{0}$
By putting $\Theta_{1}=\{(\Gamma,z_{0})\}$ and $\Theta_{2}=\{(\Lambda,x_{0}),(\Lambda_{a},\lambda_{a})\}$, we see that $(\Theta_{1},\Theta_{2},\Psi)$ is a context-free ensemble, thus the free product $(\mathfrak{C},\gamma)=(\Theta_{1}\star_{\Psi}\Theta_{2},z_{0})$ defines a context-free inverse graph called the comb digraph, see Fig. 6. Note that with the convention of the free product that we have given, the root of this inverse graph is $z_{0}$ which is also identified with $x_{0}$. In $\mathfrak{C}$ there are infinite copies of the $c$-line, we denote by $z_{i}^{(j)}$ the $i$-th vertex of the $j$-th copy of this line. Graphically, two vertices $z_{i}^{(j)}$, $z_{i}^{(k)}$ are at the same “$i$-th level”, but they belong to the $j$-th and $k$-th $c$-line, respectively. The $a$-line intersects at the vertices $z_{0}^{(j)}$.
Proposition 10.5.
The transition group $G=\mathcal{G}(\mathfrak{C})$ contains a copy of the infinite rank free abelian group $\mathbb{Z}^{\infty}$. In particular, $G\in{\bf CF-TR}$ is not a poly-context-free group.
Proof.
Consider the set $B=\{g_{n}=c^{n}ac^{-n},n\in\mathbb{Z}\}$. It is not difficult to check that for any vertex $z_{i}^{(j)}$ the generic element $g_{n}^{k}$ acts as follows:
(12) $z_{i}^{(j)}\cdot g_{n}^{k}=\begin{cases}z_{i}^{(j)},\,\,&\mbox{ if }i\neq-n\\ z_{i}^{(j+k)},\,\,&\mbox{ otherwise. }\end{cases}$
for any $j,i,k,n\in\mathbb{Z}$. Now, take $g_{n},g_{m}$ with $m\neq n$. For any vertex $z_{i}^{(j)}$ with $i\neq-n,-m$ we have
$z_{i}^{(j)}\cdot(g_{n}g_{m})=z_{i}^{(j)}=z_{i}^{(j)}\cdot(g_{m}g_{n})$
if $i=-n$, then we have
$z_{i}^{(j)}\cdot(g_{n}g_{m})=z_{i}^{(j+1)}\cdot g_{m}=z_{i}^{(j+1)}=(z_{i}^{(j)}\cdot g_{m})\cdot g_{n}=z_{i}^{(j)}\cdot(g_{m}g_{n})$
and similarly for $i=-m$. Hence, we have $g_{n}g_{m}=g_{m}g_{n}$ in $G$. Moreover, by the action described in (12) we deduce that all the elements in $B$ are distinct and each $g_{n}$ does not have torsion. Therefore, the subgroup $H=\langle B\rangle$ is an abelian subgroup of $G$. To show that $H\simeq\mathbb{Z}^{\infty}$ it is enough to prove that the set $B$ freely generate $H$, that is, there is no relation of the form
$g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}}=\mathds{1}_{G}$
where the word $g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}}$ is freely reduced. Let us assume, by contradiction, that such a relation exists. Since $H$ is commutative we may assume without loss of generality that all $k_{i}\neq 0$, and $n_{t}\neq n_{j}$ for all $t,j\in[1,\ell]$ with $t\neq j$. Now, by applying this word to $z_{-n_{1}}^{(j)}$, by (12) we would have $z_{-n_{1}}^{(j)}\cdot(g_{n_{1}}^{k_{1}}\ldots g_{\ell}^{k_{\ell}})=z_{-n_{1}}^{(j+k_{1})}\neq z_{-n_{1}}^{(j)}$ since $k_{1}\neq 0$ and $n_{1}\neq n_{j}$ for $j\neq 1$. However, this is a contradiction. Thus, since each $g_{n}$ is torsion free, we may conclude that $H\simeq\mathbb{Z}^{\infty}$. The last claim of the proposition follows by the fact that a poly-context-free group can not contain $\mathbb{Z}^{\infty}$, as it contains free abelian subgroups of rank $k$ for all $k\in\mathbb{N}$, see [9, Lemma 4.6]. ∎
Figure 7. The context-free inverse graph $\mathfrak{T}$ obtained by taking the free product of the inverse graphs in the upper part, where the gluing maps are defined by taking $\psi^{1}(z_{i})=\lambda_{a}$ for all $z_{i}\neq z_{0}$, and $\psi^{2}(x_{1})=z_{0}$.
We finally address the second point of Question 6.5 regarding the possibility for a transition group to have arbitrarily large finite groups. We prove that such examples may exists by showing a group in CF-TR having an infinite torsion subgroup. The associated graph $\mathfrak{T}$ on the alphabet $A=\{a,c,a^{-1},c^{-1}\}$ is depicted in Fig. 7. Using the notation in Subsection 7 we notice that $\mathfrak{T}$ is locally-quasi-transitive with infinite orbits by taking $n=1$. We have the following proposition.
Proposition 10.6.
The boundary subgroup $\mathrsfs{B}_{1}$ is an infinite torsion normal subgroup of $\mathcal{G}(\mathfrak{T})$ whose elements have order two. Moreover, $G/_{\mathrsfs{B}_{1}}\simeq\mathbb{Z}$.
Proof.
Let $\chi_{c}:A^{*}\to\mathbb{Z}$ be the homomorphism counting the $c$, i.e., $\chi_{c}(c)=-\chi_{c}(c^{-1})=1$ and deleting all the occurrences of $a$. Note that $\mathfrak{T}$ is locally-quasi-transitive with $\mathfrak{T}\setminus D_{1}(x_{0})=\Lambda_{1}\setminus D_{1}(y_{0})$, where $\Lambda_{1}$ is the transitive graph formed by two $c$-lines with two loops labeled by $a,a^{-1}$ everywhere. By Lemma 7.5, the language $\mathcal{O}_{1}$ is formed by words that label a circuit at every vertex of $\Lambda_{1}$. It is easy to see that this is exactly the set of words $u\in A^{*}$ with $\chi_{c}(u)=0$. Therefore, for any vertex $z_{i}^{(j)}$, $j\in\{0,1\}$ of $\mathfrak{T}$ and $u\in\mathcal{O}_{1}$ we have that either $z_{i}^{(j)}\cdot u=z_{i}^{(j)}$, or $z_{i}^{(j)}\cdot u=z_{i}^{(j+1\mod 2)}$, where the second case occurs if and only the walk $z_{i}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}z_{i}^{(j+1\mod 2)}$ passes from the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$, an odd number of times. In particular we have $z_{i}^{(j)}\cdot u^{2}=z_{i}^{(j)}$ for every vertex of $\mathfrak{T}$, whence every element of $\mathrsfs{B}_{1}$ has torsion of order two. The boundary subgroup $\mathrsfs{B}_{1}$ is infinite. Indeed, take the elements $h_{n}=c^{-n}ac^{n}\in\mathrsfs{B}_{1}$, $n\in\mathbb{Z}$. These are all distinct elements in $\mathcal{G}(\mathfrak{T})$ since for any $j\in\{0,1\}$ we have
$z_{m}^{(j)}\cdot h_{n}=\begin{cases}z_{m}^{(j+1\mod 2)}\,&\mbox{ if }m=n\\ z_{m}^{(j)}\,&\mbox{ otherwise }\end{cases}$
The last statement $G/_{\mathrsfs{B}_{1}}\simeq\mathbb{Z}$ is a consequence of $\mathcal{G}(\Lambda_{1})=\langle c\rangle\simeq\mathbb{Z}\leq\mathcal{G}(\mathfrak{T})$ and Theorem 7.6. ∎
We may explicitly characterize the word problem $\mathcal{L}(\mathfrak{T})$. We first need some definitions. For a word $u\in A^{*}$, let us denote by $u[i]$ the letter of $u$ occurring at position $i$, and by $u[:i]$ let us denote the prefix of $u$ of length $u$. We say that an occurrence of the letter $a=u[i]$, or $a^{-1}=u[i]$, is at level $m$ if $\chi_{c}(u[:i])=m$. Since there are just two vertices, namely $z_{0}^{(1)},z_{0}^{(1)}$ in which a walk may enter, it is not difficult to see that for a fixed $m$ the number of $a,a^{-1}$ at level $m$ in $u$ represents the number of times a walk starting from $z_{-m}^{(k)}$, $k\in\{0,1\}$, passes through the “bridge” given by the edges $z_{0}^{(j)}\smash{\stackrel{{\scriptstyle\text{\tiny{$x$}}}}{{\relbar\joinrel\longrightarrow}}}z_{0}^{(j+1\mod 2)}$, $x\in\{a,a^{-1}\}$. Therefore, a word $u\in A^{*}$ represents the identity in $\mathcal{G}(\mathfrak{T})$ if and only if $\chi_{c}(u)=0$ and for each $m\in\mathbb{Z}$ the number of $a,a^{-1}$ at level $m$ is even. Note that we clearly have finitely many $m$ with non-zero number of $a,a^{-1}$. We know that this language is CoCF, however, is it also poly-context-free? Let us state this in the following open problem.
Question 10.7.
Is the set of words $u\in A^{*}$ with $\chi_{c}(u)=0$ and having a number of $a,a^{-1}$ that is even at each level a poly-context-free language? This would disprove Brough Conjecture 1
Question 10.8.
Motivated by the previous constructions, it would be interesting to find examples of CF-TR groups which are not embeddable in Thompson $V$. If they existed, they would then provide counterexamples to Lehnert Conjecture.